### v14

# Technology, Institution, and Regional Growth: Evidence from Mineral Mining Industry in Industrializing Japan

Abstract
Coal extraction was an influential economic activity in interwar Japan. Initially, coal mines employed both males and females as the workforce in the pits. However, the innovation of labor-saving technologies and the renewal of traditional extraction methodology induced institutional change. This was manifested by the revision of labor regulations affecting female miners in the early 1930s. This dramatically changed the mining workplace, making skilled males the principal miners engaged in underground work. This paper investigates the impact of coal mining on regional growth and assesses how the institutional changes induced by the amended labor regulations affected its processes. By linking the mines’ location information with both registration and census-based statistics, it was found that coal mines led to remarkable population growth. The labor regulations accelerated local population growth by forcing female miners to exit the labor market and form families. The regulations prohibited female miners from risky underground work. This reduction in occupational hazard also improved early-life mortality via the mortality selection mechanism in utero.
Keywords: female labor; institution; mines; occupational hazard; regional economy; labor regulation; resource;
JEL Codes: N30; N40; N50; N70; N90;
1 Introduction
How institutions influence the economy has attracted wide attention (Hall and Jones 1999; Acemoglu et al. 2001). Institutional interventions in the labor market may impede development by preventing labor market and structural adjustments (Freeman 1993; Fernández-Villaverde 2018). International evidence indicates that strong regulation over labor markets can be associated with lower labor force participation and higher unemployment (Botero et al. 2004). [^1] Djankov et al. (2002) used a dataset on start-up firms’ entry regulations from 85 countries. They found that more robust regulation of start-up firms’ entry can be associated with higher corruption, larger unofficial economies, and worse quality of public and private goods. A different view by Nickell and Laynard (1999) argues that, in European countries, labor institutions such as labor unions are more likely to be associated with economic performance than institutional regulation. Evidence based on the domestic variations in labor regulation shows that in India, pro-worker labor regulation decreases economic activity in the formal manufacturing sector (Besley and Burgess 2004). However, a specific type of regulation, such as unemployment insurance, may improve risk-sharing and reduce the transaction cost of job searching (Acemoglu and Shimer 1999). Therefore, understanding the type of labor regulations that could hinder (or improve) economic performance is still an influential research agenda.
This paper investigates how coal mining impacted regional development during interwar Japan, and how it was influenced by the institutional change concerning female labor patterns induced by technology shocks. In prewar Japan, the coal mining sector relied on both male and female miners. However, during the interwar period there was innovation through labor-saving technologies such as coal cutters and conveyors combined with the renewal of traditional extraction methodology. Simultaneously there was revision of labor regulations affecting female miners. It dramatically changed the mining workplace from one where both female and male manual miners worked in the pits to one where skilled male miners became the principal workers. This institutional change impacted the regional economy by increasing the gender wage gap.
I found that the presence of coal mines leads to population growth, indicating regional development in the mining area. The revised labor regulations did not stagnate the local population growth but accelerated it. This was because the institutional change due to the regulations forced female miners to exit the labor market and establish families. Since the revised regulation prohibited female miners from risky work in the pits, it substantially reduced their occupational hazards. This led to declines in overall female mortality and improvements in the early-life mortality rates via the mortality selection mechanism in utero.
This paper contributes to the literature on the role of the state and institutions during economic development from the perspective of female labor regulations (Botero et al. 2004; Besley and Burgess 2004). Importantly, it is the first study investigating the impact of institutional change due to changes in the regulations governing female workers in the resource extraction sector. Coal is necessary for industrialization because wide variety of economic activities heavily rely on coal consumption. I showed that technology-induced labor regulations reduced female labor force participation and accelerated the local population growth through the family formation channel. This institutional change mitigated the occupational hazards for female miners and improved the early-life health status in the local economy. From this perspective, the findings shed new light on the role of gender-specific labor regulations in understanding the mechanisms behind regional development.
This study is relevant to the broader literature on the economic impact of resource extraction. Predictions regarding agglomeration theory imply that resource abundance can increase the demand for local labor in the mining sector and related industries, thereby causing an influx of population and agglomeration (Rosenthal and Strange 2004; Duranton and Puga 2004; Lederman and Maloney 2007). [^2] Induced agglomeration can improve efficiency in the local labor markets through, for instance, better matching and learning. This can have spillover effects in the non-mining sectors (Black et al. 2005; Bjørnland and Thorsrud 2016). See Corden and Neary (1982); Eastwood and Venables (1982); Corden (1984); van Wijnbergen (1984); Neary and van Wijnbergen (1984) for earlier theoretical studies on the impact of resource discoveries. However, growth of the resource extraction industry may stagnate the development of non-resource extraction industries by reallocating resources from the non-resource tradable sector (Sachs and Warner 1999, 2001). [^3] The wealth effect mechanism suggests that resource abundance increases the demand for commodities, thereby reallocating resources from the tradable to non-tradable sector and increasing the imports of tradables. This can reduce labor supply, leading to greater earnings in the local economy. The direct effect on the relative demand of extraction firms and workers is higher prices of personal and business services. Other political-economic factors that may disrupt the growth of the non-resource sector are rent-seeking and corruption. Caselli and Guy (2009, 2013) summarize these theoretical implications. See Papyrakis and Gerlagh (2007) for empirical evidence using United States data. There is still debate on whether proximity to coal mines was necessary for regional development during Europe’s industrialization. A recent study by Fernihough and O’Rourke (2020) found that after the mid-eighteenth century, European cities closer to coalfields had larger populations than other cities. [^4] Similarly, some economic history studies have argued for the importance of better access to coalfields in subsequent industrial development (Mathias 2001; Pollard 1981). By contrast, Mokyr (1977) observed that the distribution of coal supplies did not contribute to industrialization. My main results support the former assertion by showing that the coal mines led to population growth in the local economy. Importantly, this study contributes to the literature analyzing the mechanism behind population growth due to resource extraction. [^5] Although studies have analyzed the impact of resource abundance on population growth, the mechanisms underlying this relationship have been understudied. See for example Black et al. (2005); Michaels (2010); Fernihough and O’Rourke (2020). I demonstrate that while the population growth in the mining area initially occurred due to internal migration, the miner households formed families in response to the institutional change.
This paper also expands understanding of the relationship between mining and regional human capital accumulation. A pivotal risk of mining is pollution, which can increase the health costs in resource-abundant locations and decrease agricultural productivity (Aragón and Rud 2015; von der Goltz and Barnwal 2019). Academic consensus is that the potential health risks of mining during historical industrialization are still to be evaluated. [^6] Economic history studies have focused on the health costs of air pollution. Heblich et al. (2021) and Beach and Hanlon (2017) found that extensive coal use during historical economic growth in Britain increased infant mortality rates and decreased the average adult’s height. Relatedly, Hanlon (2020) found that coal-related pollution disturbed the growth of United Kingdom (UK)cities in the late nineteenth century. Another strand of the literature investigates the adverse impact of using lead pipes on mortality rates (Grönqvist et al. 2020). I found that while there is little evidence on the air and water pollution pathways, occupational hazards for female miners led to mortality selection in utero, which increased early-life mortality. Regulations for female miners are suggested to mitigate the mortality sorting mechanisms before births. The finding sheds light on the importance of occupational hazards in evaluating the economic activities of the resource extraction industry, which has been neglected in the literature. In addition, it provides evidence suggesting that labor regulations may improve the human capital of their offspring by improving early-life health conditions (e.g., Almond et al. 2018).
Finally, this study utilizes data constructed using census-based statistics. [^7] Most studies have focused on large-scale mines, particularly gold, and utilized location information combined with survey datasets (Aragón and Rud 2013, 2015; Kotsadam and Tolonen 2016). This enables the elimination of risks caused by selection bias in the statistical inference. This study also exploits specific geological strata given in nature as an instrumental variable. This offers robust estimates for evaluating the impact of coal mines on the regional economy. In the literature, this study provides the first evidence from a developing Asian country in the past. This significant given that earlier research has mainly focused on European countries, the US, Latin America, and African countries. [^8] See, for example, Domenech (2008); Fernihough and O’Rourke (2020) for European countries, Black et al. (2005); Michaels (2010) for the US, Sachs and Warner (1999); Aragón and Rud (2013) for Latin America, and Aragón and Rud (2015); Benshaul-Tolonen (2019) for African countries.
The remainder of the study is organized as follows. Section 2 briefly overviews the historical context. Section 3 introduces the conceptual framework for this paper. Section 4 describes the data and Section 5 summarizes empirical analysis. Section 6 concludes the study.
2 Historical Background
2.1 Coal Mining Industry
Figure 1: Coal Production in Japan from 1912 to 1940
Note: The solid line shows the coal output in giga tons (Gt). The dashed line indicates the coal price in yen per giga tons. The unit price is defined as the total price of coal output divided by the total coal output in each year.
Source: Created by the author. Data on coal output are from Ministry of International Trade and Industry (1954).
The Mining Code of 1892 allowed private companies to have mining concessions, accelerating the installation of modern technologies in mines and the development of Japan’s mining sector. Winding engines were installed in coal and copper mines in the 1890s, and the flame-smelting process was adopted in gold and silver mines in the 1900s (Ishii 1991, pp. 220–221). Before the First World War (WWI), however, the Japanese coal industry had developed as a labor-intensive export industry. Then, the outbreak of WWI increased miners’ average wage rates, reducing the coal industry’s comparative advantage. Figure 1 illustrates the development of the coal industry and shows the negative shock in the early 1920s. This motivated coal mining firms to reduce the number of miners and adopt new extraction machines to improve productivity. [^9] See Sumiya (1968, pp. 387–393) for details on the machinery adoption process. Mining firms reallocated resources to relatively productive mines to improve their average productivity. Consequently, average labor productivity in the coal mining sector increased from the 1920s, particularly in the 1930s (Okazaki 2021). [^10] Generally, during the interwar period, the mining and manufacturing industries experienced steady capital investment, particularly in production facilities and machines, and public investments (Nakamura 1971, pp. 138–143). Consequently, the relative contribution of manufacturing and mining, which provided energy and resources to manufacturing, was the largest at 42-51% among all economic sectors during this period (Minami 1994, p.90). Figure 1 shows that total coal production increased during the interwar period. [^11] As the moderate decline in the early 1930s confirmed, the impact of the depression was not substantial compared to the other countries. See Online Appendix A.1 for details.
In the early 1930s, coal accounted for approximately $55$% of total mineral output in Japan, meaning that coal extraction was an influential economic activity at that time (Mining Bureau, Ministry of Commerce and Industry 1934, p.44). The higher wage rates of miners motivated peasants to move into the extraction industry, shifting the local industrial structure from the agricultural to the production sector. An example, the average daily wage of female miners working in the pits was one and a half to two times higher than those of the female workers in the silk and spinning industries (Nishinarita 1985, pp. 84–85). Besides the original inhabitants, therefore, mining workers would include a certain proportion of migrants. This may have influenced the structure of both the marriage market and fertility in the local economy surrounding mines. However, quantitative evidence of these changes in the regional economy is still lacking in literature. [^12] Most research on mines in prewar Japan comprises business history studies that investigated the features of the management and industrial organization in coal mining companies (Ishii 1991, pp. 220-233). Tanaka (1984) provided a comprehensive review of the operations of the Japanese coal mining industry.
2.2 Mining Technology, Occupational Hazards, and Regulations of Female Miners
By the end of the 19th century, a large proportion of mines had completed mechanizing the hauling processes. This was done by installing equipment such as drainage pumps and hoisting machines. However, this did not mean that the miners’ workload related to hauling operations was reduced. Instead, such mechanization in the main shafts increased the workload of miners extracting and transporting coal around the coalface (Sumiya 1968, pp. 310–312). Moreover, the environment in the pit was generally poor. The air was polluted by dust, oxygen levels were low, and there was a risk of carbon monoxide poisoning. There was also a high fire risk from dynamite blasting and spontaneous combustion of gas and coal. Accidents caused by rockfall occurred frequently. Available statistics indicate that the average number of accidents per year in the pits between $1917$ and $1926$, causing miners’ injuries and deaths, was $142,595$, of which $61,112$ were from rockfall. Accidents caused by explosions, poisoning, and suffocation were fewer than those caused by rockfall, but they still occurred $311$ times per year on average between $1917$ and $1926$. [^13] These statistics are from Japan Mining Association (1928, p. table.6).
Notably, while miners included single male workers, many of them were married couples. This was because miners usually worked in teams of two to three (Sumiya 1968, pp. 300–308). The male skilled miner, called saki yama, extracted coal at the face; meanwhile, the female miner, named ato yama, carried that coal to the coal wagon at the gangway. She traveled through steep pits carrying bamboo baskets called sura (Sumiya 1968, p. 319). As coal is a bulky resource, the heavy work burden on pregnant women could have increased the risk of adverse pregnancy outcomes (Ahmed and Jaakkola 2007). Miners worked in pits with dirty air containing particulates, which eventually damaged their lungs and increased the risk of respiratory diseases such as tuberculosis (Sumiya 1968, pp.  410–411). [^14] A miners’ health survey found that coal miners had the highest respiratory disease prevalence among all types of miners at that time, which indicated how hard the work in the coal mines was (Japan Mining Association 1928, p. 3). Relatedly, a growing body of medical literature has revealed that particulates and radionuclides from coal increase the risk of stillbirths and premature deaths due to mothers’ respiratory diseases (Landrigan et al. 2017; Lin et al. 2013).
An important event that changed the harsh labor pattern experienced by female miners was an institutional change occurring in the early 1930s. Following the International Labor Conference after WWI, the Social Affairs Bureau held a consultation with mining companies in the 1920s to revise the Miners’ Labor Assistance Regulations of 1916. Initially, mining companies, who relied heavily on female miners for production, opposed the Social Affairs Bureau’s insistence on prohibiting women from working underground and late-night. However, the renewal of mining methods driven by new extraction technologies changed the attitudes of the mining companies in the late 1920s. The transition from the traditional room-and-pillar mining method to the longwall method was accompanied by the adoption of technologies such as coal cutters and conveyors. Compared to the room-and-pillar method, the coal extraction space of the longwall method is larger, allowing for the introduction of those machines (Nishinarita 1985, p. 91). Specifically, the conveyors brought extracted coals from pits to the gangway more efficiently than the female miners, strongly incentivizing the mining companies to prohibit underground work by female miners (Tanaka and Ogino 1977, p. 79). [^15] Online Appendix A.2 provides finer details in this mechanization process. Online Appendix A.3 provides cross-sectional evidence on the replacement of female miners by machines. Finally, the Revised Miners’ Labor Assistance Regulations of September 1928 prohibited some underground and late-night work by female miners. As the grace period for enforcement of the regulations lasted until 1933, the number of female miners gradually declined from around 1930. This change induced by mechanization was commonly observed in Japanese coal mines (Tanaka 1984, p. 392). [^16] A detailed chronology of technology adoptions in each coal mine in Chikuhō coalfield confirms this movement (Compilation Committee of the Chronology of Chikuhō Coal Mining Industry 1973).
Panel A of Table 1 demonstrates the evidence that the number of female miners in the pits reduced from 1930 to 1933 when the revised regulation was enacted. The full enforcement of the revised labor regulations did not eliminate female miners altogether because females were still allowed to work in the thin layer pits. They were primarily responsible for coal selection (sentan) and other out-of-pit labor. In the late 1930s, more than $15,000$ women still worked in the mines, representing approximately $10$% of all the miners measured. From this perspective, female miners played an important role in the coal mines throughout the interwar period. The enforcement of regulations, however, impacted the market values of female workers’ skills in the mining area. Column 1 of Panel B in Table 1 presents evidence that female miners’ relative wage rate dropped after the enforcement. While there was a decreasing trend in the relative wage rate during the mechanization process in the 1920s, the difference between 1930 and 1933 shows the largest decline ($0.66$ v. $0.48$). This reduction was not a macroeconomic trend around the mining region, given that the relative wage rate in the agricultural sector was stable during the same period (Column 2 of Panel B). Moreover, the rate for in-pit workers reduced from $0.78$ to $0.62$, whereas that for out-pit workers was relatively stable from $0.49$ to $0.45$. The mean difference before and after the enforcement year is statistically significant only in the in-pit workers’ category (bottom of Panel B). This indicates that regulating in-pit female miners did not improve the relative wage rate of the out-pit female miners. Therefore, although prohibiting female workers in the risky deep layers should mitigate the occupational hazards for female miners, it shall lower the relative wages for females in the mining area.
Table 1: Number of Coal Miners and Relative Wage Rate
Measured in the Manufacturing Censuses
Panel A: Number of miners
Female Male % share of
Year Miners Work in the pits (%) Miners Work in the pits (%) female miners
1924 65147 -- 179986 -- 27
1927 56956 71 176524 76 24
1930 32680 60 153876 76 18
1933 15695 37 133460 77 11
1936 18694 25 178816 77 9
Panel B: Relative wage rate (female/male)
(1) Mining Sector (2) Agricultural Sector
Year Overall In-pit worker Out-pit worker Daily labor Annual labor
1924 0.81 -- -- -- --
1927 0.74 0.83 0.52 0.77 0.72
1930 0.66 0.78 0.49 0.77 0.72
[…]
1936 0.45 0.67 0.46 0.77 0.70
Mean diff. 0.27 0.16 0.05 0.01 0.01
($p$-value) (0.016) (0.046) (0.124) (0.423) (0.769)
Notes: The statistics on the miners in this table are drawn from the Labor Statistics Field Survey Reports that surveyed mines employing 50 or more miners. % share of female miners in Panel A shows the percentage share of females relative to the total number of miners. The relative wage rate in Panel B is the female average daily wage divided by the male average daily wage. The overall ratios for the miners between 1927 and 1936 (column 1) are recalculated using the number of miners working inside and outside of the pits and their daily wages. The mean difference in the wage ratio around 1933 and $t$-statistic $p$-values for the mean difference are listed at the bottom of Panel B.
Sources: Statistics are from Statistics Bureau of the Cabinet (1926, pp. 5; 32), Statistics Bureau of the Cabinet (1930, pp. 6–7; 26; 30), Statistics Bureau of the Cabinet (1932b, pp. 6–7; 26; 30; 38), Statistics Bureau of the Cabinet (1936, pp. 41; 64; 68; 76), and Statistics Bureau of the Cabinet (1937, pp. 45; 78; 79). The average daily wages for the daily and annual agricultural laborers are obtained from Umemura et al. (1966, p. 107). The survey subject description is from Statistics Bureau of the Cabinet (1932a, p. 1) and Statistics Bureau of the Cabinet (1937, p. 1).
2.3 Mining Technology and Pollutions
Pollution may be another factor that could influence the regional economy. Generally, industrial pollution was not recognized as a common social problem in pre-war Japan. Consequently, there is little documentation of pollution due to coal mining. [^17] An exception is the Ashio Copper Mine Incident in the late 19th century (e.g., Notehelfer 1975). In the field of development economics, von der Goltz and Barnwal (2019) provides evidence that mining is associated with child stunting in nations that are still economically developing. However, anecdotes suggest that a few potential pathways generate pollution in coal mines.
First was the soot and smoke caused by coal burning to run steam-powered winding machines. Steam-powered winding engines were introduced early in the mechanization of the hauling process (Section 2.1) and exhausted smoke generated from burning coal in the boiler. However, these steam-powered winders were gradually replaced by electric winders during the interwar period (Japan Mining Association 1930, 1931). Thus, this channel may be unlikely to impact the regional population health. Second, presumably more relevant, is pollution from wastewater generated during mining. As coal mining became increasingly mechanized, some mines began introducing water-washing machines during coal sorting. Unfortunately, there is little historical documentation systematically describing pollution caused by wastewater from coal mines. However, coal sludge from coal selection started contaminating rivers at the time. An example is Onga river in the Chikuhō coalfield, which was called a “zenzai” (sweet red-bean soup) because the water was contaminated by the liquid waste (Chiba and Yamada 1964). Coal sludge contains several metal toxins, including mercury (Hg) and cadmium (Cd). These are known to be associated with the incidence of miscarriages and stillbirths (Amadi et al. 2017; Sergeant et al. 2022). This suggests the possibility that early-life mortality would have been higher in the coal mining regions with greater river accessibility. From the 1920s, however, some mines installed settling ponds to purify wastewater and recirculate it during coal cleaning (Advisory Council on Mining 1932, pp. 313–316). This could mitigate the pollution risks around the mines due to the wastewater.
3 Conceptual Framework
The regulation of female miners’ working conditions described in the previous section would influence the fertility in the mining area. The decline in females’ relative wage rate due to regulation can decrease the opportunity cost of housework, whereas it can also decrease their earnings in the labor market. While the former mechanism could lead to higher fertility through the substitution effect from consumption, the latter may result in lower fertility via the income effect. This section offers a brief overview of a theoretical framework predicting the relevant mechanism given the historical evidence about the coal mines in prewar Japan. To do so, I consider the collective model proposed by Siegel (2017) that illustrates the optimal fertility based on the couple household’s utility maximization. [^18] For the sake of brevity, a simplified model that abstains the utility from leisure and home production parameters is considered. This offers a materially similar interpretation of the optimal fertility choice to the original generalized model. See Browning et al. (2014) for a comprehensive review on the cooperative and non-cooperative models of the couples.
Let us consider that both males and females derive utilities from consumption ($c$) and having children ($b$) under an additively separable utility function as $u_{g}(c_{g},b)=\log c_{g}+\tau\log b$, where $g\in\{m,f\}$ indicates a spouse. Raising children requires a home production $y(b)=\pi b$ with $\pi>0$, which is accomplished using labor-saving technologies available in the markets ($s$) and parents’ total home labor input ($\Omega$) as: $y(b)=s^{\gamma}\Omega^{1-\gamma}$. $\gamma$ represents the availability of childcare services in the economy ($10\leq\gamma<$ 1). The parents’ total home labor input is specified in the constant elasticity of the substitution function as: $\Omega=\left(\kappa_{m}h_{m}^{1-\rho}+\kappa_{f}h_{f}^{1-\rho}\right)^{\frac{1}{1-\rho}}$, where $\kappa_{g}$ is the productivity of housework including childcare, $h_{g}\geq 0$ is the home labor input, and $\rho>0$ is the inverse of the elasticity of substitution. The labor income is derived as the product between wage $w_{g}$ and labor input $n_{g}\geq 0$, which is used for their consumption and for purchasing home labor-saving services.
The representative household maximizes the sum of utilities $\theta u_{m}(c_{m},b)+(1-\theta)u_{f}(c_{f},b)$ subject to budget constraints $c_{m}+c_{f}+s=w_{m}n_{m}+w_{f}n_{f}$ and time constraint $n_{g}+h_{g}=1~{}\text{for}~{}g\in\{m,f\}$. The optimal fertility choice ($b^{*}$) is then given in a closed-form function. One can confirm that the sign of the derivative of optimal fertility with respect to the wage of female ($\frac{\partial b^{*}}{\partial w_{f}}$) is negative under the following condition:
$\footnotesize{\begin{split}\gamma\kappa_{f}^{\frac{1}{\rho}}w_{f}^{\frac{\rho-1}{\rho}}+w_{m}\left[\left(\frac{\kappa_{m}}{w_{m}}\right)^{\frac{1}{\rho}}-(1-\gamma)\left(\frac{\kappa_{f}}{w_{f}}\right)^{\frac{1}{\rho}}\right]<0.\end{split}}$ (1)
Online Appendix B provides details of the derivation. Under a paternalistic social norm stipulated in the Old Civil Code, the head’s housework productivity, particularly childrearing, was negligible ($\kappa_{m}=0$). It follows that condition 1 implies $\frac{w_{f}}{w_{m}}<\frac{1-\gamma}{\gamma}$ for my empirical setting. Since the marketization of childcare ($\gamma$) was considerably low at that time, the optimal fertility was also negatively associated with the relative wage ($w_{f}/w_{m}$). [^19] As mechanization progressed, miners were required to accumulate knowledge and experience in machinery operation. Management, therefore, had an incentive to encourage miners to form families and continue employment over a long term (Ogino 1993). Thus, many coal mines were equipped with nurseries and other welfare facilities from the 1920s (Ministry of Commerce and Industry 1926). However, evidence shows that these nurseries required fees and had covered only a small part of their childcare requirements. See Online Appendix B for details. This result suggests the following two predictions:
  1. Prediction 1: Before female labor regulation, the mining area could have lower fertility, given the higher relative wages in the mining sector.
  2. Prediction 2: After female labor regulation, the mining area could have recovered fertility in response to the reduction in the relative wages.
The short-run population dynamics in a local economy depend on natural and social changes. [^20] From year $t$ to $t+1$, this geometric population growth model is specified as $\Psi_{t+1}-\Psi_{t}=(\Upsilon_{t}-\Theta_{t})+\Pi_{t}$, where $\Psi$, $\Upsilon$, $\Theta$, and $\Pi$ indicate the number of people, live births, deaths, and net migrants, respectively. In the former change, the lower fertility under the substitution effect before the regulation would not be a major factor in population growth in the mining area. In turn, the fertility gain after the regulation could increase population growth. Mortality term in the natural change is given outside of the household’s fertility choice in the short-run. The regulation of female miners reduced their occupational hazards, which could improve female mortality. The improvements in maternal health could also mitigate prenatal and neonatal deaths (Section 2.3). Both ameliorations would work to reduce population losses in the mining area. The influence of social change on local population growth is uniquely determined when the sign of natural change is determined. Moreover, social change is positive in the mining area because the regional wage gap would encourage the influx of workers from surrounding agrarian areas into the mining area. For female workers, the wage reduction via the regulation could reduce the attractiveness of coal mining labor. However, it may not offset the benefits for couples moving into the mining area because male miners’ wages were not reduced by the regulation. [^21] Thus, since female workers migrated with their husbands, the migration of couples is less affected by the regulation than that of single workers. To summarize, while the mining area had an influx of workers throughout the period under consideration, the fertility and mortality in the mining area could be affected by the enforcement of the revised regulation of 1933.
4 Data
This study created a unique dataset of the demographics across mining areas in the Japanese archipelago. As mining is a localized economic activity, a smaller lattice dataset is preferred to identify the impact of mines. A set of official census-based municipal-level statistics published in the 1930s was collected and digitized.
4.1 Mine Deposit
I use official reports named Zenkoku kōjyō kōzan meibo (lists of factories and mines) published by Association for Harmonious Cooperation (1932, 1937) (hereafter, called the AHC), which document coal mine locations measured in October 1931 and October 1936. The AHC listed all mines employing $50$ miners and more. Thus, the main target of this study is the average impact of small- to large-scale coal mines and not the micro-scale collieries employing fewer than $50$ mimers. [^22] Generally, several coal mines are clustered in a coal mining municipality. For example, 85 (108) out of 93 (115) municipalities with mines had one to four mines in the 1930 (1935) sample. In this sense, this study estimates the average impact of these general coal mining municipalities. In Online Appendix D.11, I confirmed that the main findings are not sensitive to a small number of municipalities with several mines. Despite this, the AHC is still comprehensive as it includes information on small-scale mines, which have been neglected in the literature.
The treatment group is defined as the municipalities located within 0–5 km of the centroid of a municipality with mines. [^23] The study used an official shapefile provided by the Ministry of Land, Infrastructure, Transport and Tourism for geocoding. See Online Appendix C.1 for the details. A 5 km threshold is fundamentally plausible because the mean value of the distance to the nearest neighborhood municipality is approximately 4 km. In addition, the estimated effects disappeared in the statistical sense outside the 5 km range, which means that the potential spillover effect is captured under this threshold. Online Appendix D.6 summarizes finer details of the validity of the threshold. For the control group, the threshold is set as 5–30 km from the centroid. [^24] The definition of both groups is conservative compared with that in the literature, which uses 10–20 km and 100–200 km as the thresholds of the treatment and control groups, respectively (Wilson 2012; Aragón and Rud 2015; Kotsadam and Tolonen 2016; Benshaul-Tolonen 2019). This is because Japan has a smaller and thinner archipelago than countries studied in the literature, such as South Africa (Online Appendix Figure 2). In Online Appendix D.6, further evidence is provided that 5–30 km is a plausible distance for the control group given that the impacts of the coal mines are concentrated within 5 km distance from the mines.
(a) 1931
(b) 1936
Figure 2: Spatial Distribution of the Treated and Controlled Municipalities
Notes: The white circles indicate the location of municipalities with coal mines. The treatment group highlighted in red includes the municipalities within 5 kilometers of a mine. The control group highlighted in pink includes the municipalities between 5 and 30 kilometers from a mine. The excluded municipalities are shown as empty lattices in the figures.
Source: Created by the author.
(a) 1931
(b) 1936
Figure 3: Location of Coal Mines around Chikuhō Coalfield in Kyūshū
Notes: The white circles indicate the location of municipalities with coal mines around Chikuhō coalfield (Figure 2a) in Kyūshū region. Figure 3a shows the coal mines in Chikuhō, Kasuya, and Miike coalfields in Fukuoka prefecture, Karatsu coalfield in Saga prefecture, and Miike coalfield in Kumamoto prefecture. Amakusa coalfield in Kumamoto prefecture is added in Figure 3b (on the lower left of the figure). Ōita prefecture does not have coal mines but is shown in the figures to explain the border of the sample. Treatment (control) group includes municipalities within 5 (between 5 and 30) km from a mine. The excluded municipalities are shown as empty lattices in the figures.
Source: Created by the author.
Figure 2 indicates the coal mine locations from the AHC and the corresponding treatment and control groups by year. Most coal mines are located in three specific areas in the southernmost, central, and northeastern regions. These are representative coalfields called Chikuhō, Jyōban, and Ishikari-Kushiro Coalfields, respectively. Notably, a comparison of Figures 2a and 2b shows that these agglomerations were stable over time. For a closer look, Figure 3 provides an example map for the most representative coalfield called Chikuhō Coalfield in Kyūshū, Japan’s southernmost island (indicated in Figure 2a). One can observe that the distribution of mines has changed little, with only a few mines either opened or closed between 1931 and 1936. This feature of coal mines makes it difficult to use within-variation in the spatiotemporal distribution of the mines for identification. Online Appendix D.8 discusses this point in detail. [^25] In short, it is not feasible to implement a difference-in-differences estimator for the full sample because of the lack of within-variation (i.e., the number of municipalities that experienced the opening/closing of the coal mines). Nevertheless, in Online Appendix D.8, the results from the two-group by two-period difference-in-differences setting for a small subsample including Ogawa and Miyoshiyama coal mines will be demonstrate (Figure 2b), which provide the results reasonable to the findings.
Finally, the generation of the analytical samples will be explained. From $11,151$ total municipalities, the municipalities within a $30$ km radius from the centroid of the municipalities with mines are first retained. The intersection of the different types of mines are then excluded, such as municipalities within $5$ km from a gold, silver, and copper (GSC) mine. Online Appendix Table C.3 presents the summary statistics of the treatment variables. The number of municipalities in the analytical sample based on the 1931 and 1936 AHC were $1,140$ and $1,364$, respectively. Meanwhile, the proportion of treated municipalities was approximately $15$% and $13$%, respectively, for each sample. [^26] The number of municipalities with coal mines in each analytical sample is 93 and 115, respectively. The slight reduction in the share of the treatment group is due to the increase in the number of enclaves, which is explained above. A mine opening in the enclave area increases the number of controlled (i.e., surrounding) municipalities, which leads to a decrease in the relative share of the treatment group. Examples are the Ogawa and Miyoshiyama coal mines in Tōhoku region (Figure 2b).
4.2 Geological Stratum
The spatial distribution of a specific geological stratum created in the Cenozoic era, which includes some carboniferous ages, is used as an instrumental variable (IV) for the location of coal mines (Section 5.1). Data on the geological strata are obtained from the official database of the Ministry of Land, Infrastructure, Transport and Tourism, which includes roughly nine thousand stratum points. Each municipality in the analytical sample was matched with the nearest stratum point to identify the municipalities with the relevant stratum. Online Appendix C.2 provides finer details of the definition and an example of a geological columnar section of a representative coal mine to show the relevance of the IV. Online Appendix Table C.3 presents the summary statistics of the indicator variables for the stratum.
4.3 Dependent Variables
The primary dependent variable, population, was obtained by digitizing the municipal-level statistics documented in the 1930 and 1935 Population Censuses (Statistics Bureau of the Cabinet 1935, 1939).
To investigate the natural change mechanism behind the local population growth due to mining, the Municipal Vital Statistics of 1930 and 1935 (Statistics Bureau of the Cabinet 1932c, 1938) were digitized. They documented the total number of marriages, live births, and deaths in all municipalities in the survey years. Four dependent variables were then considered: crude marriage rate, crude birth rate, marital fertility rate, and gender-specific death rate. [^27] The conventional definitions of these variables were followed. The crude marriage and birth rates are the number of marriages and live births per $1,000$ people, respectively. Marital fertility is defined as the number of live births per $1,000$ married females. Although the number of married females in each age is unavailable, the number of married females is similar to the number of households in each municipality, meaning that the marital fertility definition herein is useful in representing the couple’s fertility decision. In other words, the results from this marital fertility definition shall not be sensitive to the existence of the elder married women in the households. Although marital fertility is considered to capture the couple’s fertility decision better, the results can be materially similar between crude and marital fertilities because out-of-wedlock births have been historically rare in Japan. The crude death rate is the number of deaths per $1,000$ people. These are used to test the fertility responses of the miners and the changes in mortality due to mining activities and regulations. To further analyze the occupational hazard channel, the infant mortality rate and fetal death rate were considered. [^28] The infant mortality rate is the number of infant deaths per $1,000$ live births. The number of censored observations in infant mortality is negligible. The fetal death rate is defined as the number of fetal deaths per $1,000$ births. The censored observations in the fetal death rates are less than $11$% in both census years. The estimates from the Tobit estimator (Tobin 1958) are materially similar to those from the OLSE (not reported). This confirms the evidence that censoring is not material in the empirical setting. These rates allow for testing the impacts of the occupational hazards for female miners on the mortality selection before birth under the biological framework (Section 5.4.1). The mortality rate of children aged 1–4 was also considered to assess the potential impacts of air pollution. [^29] The child mortality rate is defined as the number of deaths of children aged 1–4 in 1933 (or 1938) per $1,000$ children aged 1–5 in 1930. Note that the number of children aged 1–5 from 1930 is used as it is only documented in the 1930 population census. Given that child deaths, especially at age 5, were rarer than infant deaths at that time, the mismatch in the age bins between numerator and denominator should not be a critical problem. The difference in the survey points does add noise to the estimation of the standard error. However, the coefficient estimate is sufficiently close to zero and thus, should not be a practical issue in this setting (Table 5). To obtain both variables, the 1933 and 1938 reports documenting municipal-level statistics on births and infant deaths published by the Aiikukai and Social Welfare Bureau were digitized.
Comprehensive statistics on migrants were unavailable in the prewar period. As described earlier, however, once the contribution of natural change to population growth is known, the sign of social change in the geometric population growth model is uniquely determined. Considering this, it is prefereable to assess the extent the influx of single male miners disturbed the gender balance, average household size, and how it changed after the regulation was enforced. The data on the sex ratio and average household size are obtained from the official reports of the 1930 and 1935 Population Censuses. [^30] The sex ratio is the number of males divided by the number of females. The average household size is the number of people in households divided by the number of households.
Overall, these variables are expected to represent the characteristics of local demographics (Angrist 2002; Abramitzky et al. 2011). Importantly, the census documents the total number of people in all municipalities in the census years. Similarly, the vital statistics are corrected under a national registration system that covers all the births and deaths in a certain year. [^31] The potential imprecision of birth data in prewar Japan was mostly eliminated by the 1920s (Drixler 2016). This comprehensiveness helps avoid sample selection issues in statistical inference. Online Appendix Table C.4 presents the summary statistics on these variables by treatment status and census year. The mean differences are statistically significant for most variables, implying that systematic demographic changes may have occurred around the mining extraction area.
4.4 Control Variables
As coal is a bulky mineral, railways were the primary transportation of the extracted coal. Firms may have preferred setting extraction points closer to railway stations to reduce transportation costs (Sumiya 1968, pp. 440–441). Further, the local population size could be larger as one moves closer to stations. Therefore, controlling for the distance between each municipality and the nearest-neighboring station is preferable to deal with the potential endogeneity issue. The location information of all stations obtained from the official dataset provided by the Ministry of Land, Infrastructure, Transport and Tourism was used to compute the nearest-neighbor distance to stations in 1931 and 1936.
Although railways were primarily used to transport coal because coal mines were usually inland, marine transportation was also used for secondary logistics (Online Appendix Figure 2). Hence, similar to railway transportation, the distance to the nearest neighboring seaport is included as a variable to control for accessibility to marine transportation.
Accessibility to rivers is also considered. Not all mining areas have large rivers suitable for water transportation, meaning that rivers did not dominate the locations of coal mines. However, several mines used river transport as primary logistics until the development of railways before the late 1920s (Chiba and Yamada 1964). This implies that the spatial distribution of the river could have influenced the location of mines in such mining areas. To control for this possibility, the distance to the nearest neighboring river is included as a control variable.
Finally, the potential influence of the elevation is taken into account. The mine’s location might have been correlated with the elevation of municipalities because terrain may relate to the locational fundamentals such as the climatic conditions and fixed cost of transportation. The grid cell elevation data provided by the Ministry of Land, Infrastructure, Transport, and Tourism was used to systematically calculate the average elevation in each municipality.
As explained below, including these control variables has little influence on the estimates after conditioning on the city and town fixed effects. To be conservative, however, it is preferable to include all the controls to ensure the orthogonality condition of the instrumental variable. Online Appendices C.4.1C.4.4 summarizes the details of the data on railways, seaports, rivers, and elevation. Online Appendix Table C.5 presents the summary statistics on these control variables. [^32] Online Appendix Figures C.5C.6, and C.7 illustrate the locations of the railway stations, seaports, and rivers in the Japanese archipelago, respectively. Online Appendix Figure C.8 shows the spatial distribution of the elevation. All the controls are log-transformed to better model the skewness of the distances and elevations in nature, albeit this transformation does not influence the results.
5 Empirical Analysis
5.1 Estimation Strategy
The random nature of mineral deposits is leveraged to identify the effects of mines. The linear regression model is as follows:
$\footnotesize{\begin{split}y_{i}=\alpha+\beta\text{{MineDeposit}}_{i}+\mathbf{x}_{i}^{\prime}\bm{\gamma}+e_{i},\end{split}}$ (2)
where $i$ indexes municipalities, $y$ is the outcome variable, MineDeposit is an indicator variable that equals one for municipalities within 5 km from a mine, $\mathbf{x}$ is a vector of control variables, and $e$ is a random error term. As the placement of a mine is determined by a geological anomaly, MineDeposit is random in nature (Benshaul-Tolonen 2019, p.1568). However, the rest of the variation may be correlated with the local variation in infrastructure. Thus, if a mineral deposit is found between a village and a city, the mining firm may choose the city as its main mining point because the city is likely to have better infrastructure than the village. Then, MineDeposit can be positively correlated with the error term, leading to a positive (negative) omitted variable bias in the estimate of $\beta$ if the placement of the city is positively (negatively) correlated with the outcome variable.333333 As an example, consider a simplified projection of $y$ on MineDeposit and City, $y=\alpha+\beta\text{{MineDeposit}}+\gamma\textit{City}+e$. When the municipality type (City) is omitted, the linear projection coefficient can be written as $\beta^{*}=\beta+\Xi\gamma$, where $\Xi=(E[\text{{MineDeposit}}^{2}]^{-1})(E[\text{{MineDeposit}}\text{{City}})$. As explained, MineDeposit and City may be positively correlated such that $\Xi>0$. In addition, when $y$ is the population, conditional on mine deposits, the city has a greater number of people than town/villages ($\gamma>0$). This implies that $\beta^{*}=\beta+\Xi\gamma>\beta$. To deal with this systematic bias in the ordinary least squares, two indicator variables are included that equal one for cities and towns, respectively. As explained in Section 4.4, the distances to the nearest neighboring railway station, seaport, and river and the elevation are included to explicitly control for transportation accessibility. [^34] The differences in the estimates from both specifications, including and excluding the control variables, indicate how much this omitted variable mechanism influences the results. Thus, the randomness of the primary exposure variable can be partially assessed (MineDeposit). As demonstrated later, the results from the specification including the city and town fixed effects are materially similar to those from the specifications. This includes the control variables, thereby supporting the randomness of the exposure variable. To be conservative, the specification including the control variables is preferred.
A potential threat in the identification may be measurement errors in time-dimensional assignments. As explained in Section 4.1, mine data were surveyed in October 1931 (1936), whereas the Population Census was conducted in October 1930 (1935). Generally, this lag concerning the censuses should not lead to a strong attenuation issue because the increments in the number of mining municipalities per year were considerably limited (Section 4.1). In turn, the data on fetal death and infant mortality were obtained from the 1933 (1938) Vital Statistics. The lags in the matching allow the consideration of the exposure durations. Thus, fetus miscarriages in 1933 (1938) were in utero conceived in 1932 (1937), whereas their mothers should have been exposed to any shocks at the time of conception in 1932 (1937). The same argument can be applied to infant mortality. Infants who died in 1933 (1938) were born at the beginning of 1932 (1937) at the earliest. Therefore, they should have been in utero in 1931 (1936). Consequently, the mining data that list the mines in October 1931 and 1936 may be reasonably matched with the health data of 1933 and 1938, respectively. However, one must be careful as there is still a one-year lag in matching with both the census and vital statistics datasets. This may lead to attenuation bias due to miss assignments in the time dimension, because the number of coal mines should be increased during that year. [^35] Note that this type of measurement error never overstates the estimates but causes attenuation. Thus, if a few treated (untreated) municipalities were regarded as the untreated (treated) municipalities, the impacts of mines shall be discounted. See Online Appendix D.1 for a brief explanation of this mechanism.
To deal with such potential issues, the IV estimator is considered using the exogenous variation in the geological stratum. In this approach, equation  2 is regarded as a structural form equation because the least-squares estimator ($\hat{\beta}$) is assumed to be attenuated by the measurement error in the exposure variable. The reduced-form equation for MineDeposit is designed as follows:
$\footnotesize{\begin{split}\textit{MineDeposit}_{i}=\kappa+\tau\textit{Stratum}_{i}+\mathbf{x}^{\prime}_{i}\bm{\mu}+\epsilon_{i},\end{split}}$ (3)
where Stratum is a binary IV that equals one for municipalities with sedimentary rock created during the Cenozoic era, and $\epsilon$ is a random error term. The IV is plausibly excluded from the structural equation because the location of mines is essentially dominated by the distribution of geological stratum, which is exogenously given in nature and unobservable by people in the prewar period. In addition, the relevance condition holds because the location of coalfields is basically determined by the distribution of the stratum created in the Cenozoic era, which includes the Carboniferous period when the strata containing coal were created (Tanaka 1984; Fernihough and O’Rourke 2020). Online Appendix C.2 summarizes the geological strata variable in finer detail and shows that the location of coal mines is determined by specific strata created in this era. Online Appendix D.2 discusses the validity of identification assumptions in more rigorous way.
The heteroskedasticity-consistent covariance matrix estimator is used as a baseline estimator. [^36] The results are materially similar between the Eicker-White type covariance matrix estimator suggested by Hinkley (1977) and the HC2 estimator proposed by Horn et al. (1975). For the reduced form regressions, to be conservative, the HC2 estimator is used given that the treatment variable is relatively sparse (Section 4.1). For the sensitivity check, the standard errors clustered at the county level based on the cluster-robust covariance matrix estimator (Arellano 1987) are used to determine the influence of the potential influences of the local-scale spatial correlations. Online Appendix D.3 shows that the results are materially similar under both variance estimators. Therefore, the potential spatial correlations were negligible in this empirical setting.
5.2 Main Results
Table 2: Coal Mining and Regional Development:
Local Population Growth in the Mining Area
Panel A: 1930 Census
Dependent Variable: ln(Population)
(1) (2) (3) (4)
MineDeposit 1.3861.386***1.386 ∗ ∗ ∗ 0.8290.829***0.829 ∗ ∗ ∗ 0.6820.682***0.682 ∗ ∗ ∗ 0.4580.458***0.458 ∗ ∗ ∗
$(0.291)$ $(0.197)$ $(0.197)$ $(0.058)$
City and Town FEs No Yes Yes Yes
Railway accessibility No No Yes Yes
Port accessibility No No Yes Yes
River accessibility No No Yes Yes
Elevation No No Yes Yes
Observations 1,140 1,140 1,140 1,140
Estimator IV/Wald IV IV OLS
First-stage $F$-statistic $43.39$ $37.70$ $36.05$-- -
Mean of the DV $8.29$ $8.29$ $8.29$ $8.29$
Std. Dev of the DV $0.74$ $0.74$ $0.74$ $0.74$
[…]
       Geometric Mean for the TG/CG $7,172/3,626$
Panel B: 1935 Census
Dependent Variable: ln(Population)
(1) (2) (3) (4)
MineDeposit 1.4011.401***1.401 ∗ ∗ ∗ 0.8700.870***0.870 ∗ ∗ ∗ 0.7220.722***0.722 ∗ ∗ ∗ 0.4690.469***0.469 ∗ ∗ ∗
$(0.278)$ $(0.188)$ $(0.181)$ $(0.056)$
City and Town FEs No Yes Yes Yes
Railway accessibility No No Yes Yes
Port accessibility No No Yes Yes
River accessibility No No Yes Yes
Elevation No No Yes Yes
Observations 1,364 1,364 1,364 1,364
Estimator IV/Wald IV IV OLS
First-stage $F$-statistic $48.87$ $44.11$ $44.95$-- -
Sample Mean of the DV $8.27$ $8.27$ $8.27$ $8.27$
Std. Dev of the DV $0.74$ $0.74$ $0.74$ $0.74$
[…]
       Geometric Mean for the TG/CG $7,708/3,744$
***, **, and * represent statistical significance at the 1%, 5%, and 10% levels, respectively. Standard errors based on the heteroskedasticity-robust covariance matrix estimator are reported in parentheses.
Notes: Panels A and B present the results for the 1930 and 1935 samples, respectively. Columns 1–3 show the estimates based on the IV estimator for the system of equations 2 and 3. Column 4 shows the estimates from the OLS estimator for equation 2. Geometric Mean in the TG and CG show the $\text{exp}(\hat{\alpha}+\hat{\beta})$ and $\text{exp}(\hat{\alpha})$ from the estimated system, respectively. Panel A of Online Appendix Table C.4 shows the summary statistics for the population (before log-transformed) by treatment and control groups.
Panel A of Table 2 presents the main results for the population in 1930. The estimate from a simple IV estimation is shown in column 1.373737The first-stage $F$-statistics reported in the same panel exceed the rule-of-thumb threshold value proposed by Staiger and Stock (1997), say 10, in all regressions. This shows that the necessary rank condition is satisfied. Online Appendix D.3 confirms that the first-stage $F$-statistics never break this threshold when a conservative variance estimator such as the cluster-robust variance-covariance matrix estimator is used. Then I include the city and town fixed effects (column 2), and both the fixed effects and control variables (column 3). The estimated coefficient decreases from $1.386$ to $0.829$ when the fixed effects are included. This suggests that, as expected, the location of the coal mines may be positively correlated with the local infrastructure. Including additional control variables on transportation accessibility has a moderate influence on the estimate ($0.682$). This supports the evidence that, after conditioning on the fixed effects, there are few influential unobservables that are potentially correlated with the location of mines. Panel B confirms the similar results for the population in 1935 in the same column layouts. The estimated coefficient from the specification including the full set of controls in column 3 is $0.723$, which is larger than that for the 1930 sample.
The estimated magnitude is economically meaningful. The estimates from the preferred specification of column (3) indicate that coal mines increased the local population by $98$% [^38] Statistically, it is more precise to state that the geometric mean of population in the treatment group with mines was approximately 98% (i.e., $\text{exp}(0.682)=\text{exp}(\hat{\alpha}+0.682)/\text{exp}(\hat{\alpha})\approx 1.98$) greater than that of the control group. However, simple interpretations are used throughout this study to avoid redundancy. in 1930 (panel A) and $106$% in 1935 (panel B). The average number of miners per coal mine increased just less than $100$ during this period ($1,567$ in 1931 and $1,658$ in 1936). [^39] This is because while the number of male miners increased, simultaneously the number of female miners decreased due to the revision of the regulation (Table 1). From this perspective, it is surprising that the magnitude is estimated to be greater in the 1935 sample. The mechanism behind this trend is assessed in the next subsections.
Column 4 presents the estimates based on the reduced-form assumption to determine whether the IV approach works as expected. Even after conditioning on the control variables, the estimates based on the IV approach in column 3 are greater than those in column 4. This implies that the IV estimation strategy deals with systematic attenuation due to measurement errors.
5.3 Fertility Responses
Table 3: Assessing Mechanisms: Marriage and Fertility Responses
Pre-regulation Post-regulation
Data: 1930 Census Data: 1935 Census
(1) Crude (2) Crude (3) Marital (4) Crude (5) Crude (6) Marital
    Marriage     Fertility     Fertility     Marriage     Fertility     Fertility
MineDeposit 2.732-2.732***- 2.732 ∗ ∗ ∗ 6.815-6.815***- 6.815 ∗ ∗ ∗ 36.792-36.792***- 36.792 ∗ ∗ ∗ 1.482-1.482**- 1.482 ∗ ∗ $-1.578$ $-9.734$
$(0.794)$ $(1.944)$ $(10.343)$ $(0.649)$ $(1.509)$ $(8.169)$
City and Town FEs Yes Yes Yes Yes Yes Yes
Railway accessibility Yes Yes Yes Yes Yes Yes
Port accessibility Yes Yes Yes Yes Yes Yes
River accessibility Yes Yes Yes Yes Yes Yes
Elevation Yes Yes Yes Yes Yes Yes
Observations 1,140 1,140 1,140 1,364 1,364 1,364
Estimator IV IV IV IV IV IV
First-stage $F$-statistic $36.05$ $36.05$ $36.05$ $44.95$ $44.95$ $44.95$
Sample Mean of the DV $8.59$ $34.06$ $174.14$ $9.04$ $34.22$ $177.50$
Std. Dev. of the DV $2.28$ $5.52$ $29.69$ $2.41$ $5.47$ $29.75$
***, **, and * represent statistical significance at the 1%, 5%, and 10% levels, respectively. Standard errors based on the heteroskedasticity-robust covariance matrix estimator are reported in parentheses.
Notes: Columns 1–3 and 4–6 present the results for the 1930 and 1935 samples, respectively. Crude marriage, crude fertility, and marital fertility indicate the crude marriage rate, crude birth rate, and marital fertility rate, respectively Panel A of Online Appendix Table C.4 shows the summary statistics on these variables by treatment and control groups.
The natural growth channel is testable using the birth and death statistics comprehensively measured in the vital statistics. If the population grows naturally from family planning, municipalities with mines should have had higher marriage and fertility rates. If the substitution effect were dominant, fertility rates in the mining area would be lower than those in the surrounding agrarian area (Section 3). Table 3 presents the results for the crude marriage and fertility rates from the same estimation strategy used in Table 2. Columns 1-3 and 4-6 show the results for the 1930 and 1935 samples, respectively. [^40] The magnitudes for different years are compared to discuss the potential mechanism behind the growth of the estimated magnitude of the coal mines on the local population. Note again that although several coal mines were opened by 1935, most already existed in 1930. Thus, the cross-sectional results from the different years are comparable because the sample composition is similar over time. All the regressions include the fixed effects as well as the control variables.
Column 1 indicates that the estimate for the crude marriage rate is negative and statistically significant. This may reflect that the influx of single male miners could be associated with lower marriage rates among the entire local population. The estimate for the crude birth rate is also negative and statistically significant (column 2). This is consistent with the lower marriage rate in the mining area and the theoretical prediction that a relatively higher wage rate for female miners before the enforcement of regulation decreased the fertility rate. To disentangle both pathways, the marital fertility rate was used as the dependent variable in column 3 to rule out the attenuation from the greater proportion of single males in the mining area. The estimated coefficient is statistically significantly negative. The estimate ($-36.79$) is more than one standard deviation of the dependent variable, suggesting an economically meaningful effect.
The results for the 1935 sample listed in columns 4–6 show a clear change after the enforcement of the female labor regulation. While column (4) shows that the estimate for the marriage rate is still statistically significantly negative, its magnitude is roughly half of the estimate for the 1930 sample. Similarly, the estimate for the crude fertility rate is roughly a quarter (column 5). Importantly, the estimate for marital fertility in column (6) shows the largest magnitude reduction from the estimate in column 3. It follows that the estimates for the fertility rates are no longer statistically significant.
The results show that marriage and fertility rates have reverted to their natural trends after the regulation was enforced. This aligns with the prediction described in Section 4, supporting the evidence that the declines in female relative wages had shifted women to take on the role of family building. [^41] The optimal ratio of home labor is given by $\frac{h_{m}}{h_{f}}=\left(\frac{\kappa_{m}}{\kappa_{f}}\frac{w_{f}}{w_{m}}\right)^{\frac{1}{\rho}}$ (Online Appendix B). The labor force participation rate is only available in the 1930 Population Census. Online Appendix D.4 confirms that the mining area had lower female labor force participation than the surrounding agrarian area, with a structural shift from the agricultural to the mining sector. This is consistent with the fact that the mining area has lower female relative wages than the agrarian area (Section 2.2), supporting the labor supply mechanism behind the fertility choice. Although the data on the labor force participation rate are unavailable in the 1935 Population Census, the time-series statistics also confirm that the revision of the regulations decreased the number of female miners (Section 2.2). Under limited substitutions with the husband’s home labor and purchased services, the increase in the gender wage gap pushed the females out of the local labor market and shifted them to domestic work. This forced women to take on the role of building a family.
5.4 Mortality Changes
Table 4: Testing the Mechanisms: Mortality Changes
Pre-regulation Post-regulation
Data: 1930 Census Data: 1935 Census
(1) Male (2) Female (3) Male (4) Female
MineDeposit $0.698$ $-1.545$ $-0.578$ 2.010-2.010*- 2.010 ∗
$(1.448)$ $(1.388)$ $(1.173)$ $(1.074)$
City and Town FEs Yes Yes Yes Yes
Railway accessibility Yes Yes Yes Yes
Port accessibility Yes Yes Yes Yes
River accessibility Yes Yes Yes Yes
Elevation Yes Yes Yes Yes
Observations 1,140 1,140 1,364 1,364
Estimator IV IV IV IV
First-stage $F$-statistic $36.05$ $36.05$ $44.95$ $44.95$
Sample Mean of the DV $20.06$ $18.59$ $18.36$ $16.97$
Std. Dev. of the DV $4.48$ $4.33$ $4.12$ $4.21$
***, **, and * represent statistical significance at the 1%, 5%, and 10% levels, respectively. Standard errors based on the heteroskedasticity-robust covariance matrix estimator are reported in parentheses.
Notes: Columns 1–2 and 3–4 show the results for the 1930 and 1935 samples, respectively. Columns 1 and 3 show the results for the male death rate, whereas columns 2 and 4 show the results for the female death rate. Panel A of Online Appendix Table C.4 summarizes the summary statistics of the mortality rates by treatment and control groups.
Table 4 presents the results for the mortality channel by gender. [^42] The result presented in this section is unchanged if the log-transformed rates is used instead of the raw rates (not reported). Columns 1–2 show the results for the 1930 census, whereas columns 3–4 show those for the 1935 sample. As shown, the estimates for the former sample are close to zero and statistically insignificant. This suggests that the mining area has a mortality rate similar to that of the surrounding area. There are a few possible interpretations of this. One possibility is that the heavy work burden in the surrounding agrarian area had balanced the mortality risk. If this channel is more relevant, the female mortality rate in the mining area could decrease after the enforcement of regulation because it does not improve the health status of the agrarian workers but improves only the health status of the female miners. Another explanation is that the higher consumption levels offset the greater risk of miners’ mortality by improving overall health status in the mining area. If this channel is closer to reality, the enforcement should not improve mortality in the mining area because it may work to reduce the overall earnings of mining households.
Turning to look at columns 3–4, the estimated coefficient for the female mortality rate alone is negative and weakly statistically significant. The estimate is $-2.010$, roughly half of the standard deviation of female mortality in 1935. Thus, this may support the former occupational hazard channel. The institutional change from the enforcement of regulation liberated women from dangerous underground work, which may have improved the overall health status of females.
5.4.1 Mortality Selection before Birth: Testing the Impacts of the Occupational Hazards and Pollution on Early-life Mortality
The result for overall mortality suggests that the regulation mitigated the female miners’ occupational hazards. To illustrate a more comprehensive view of the improvements in local mortality risk, the extent to which occupational hazards influenced the reproductive health of female miners and whether institutional change improved such hazards are assessed. This is done using the early-life mortality measured in the registration-based vital statistics. The potential influence of mining pollution in the mining area is also assessed.
Theoretical Framework
First of all, a theoretical framework is illustrated for mortality selection in utero to interpret the quantitative evidence from the vital statistics data. The biological sorting mechanism in utero suggested by the Trivers-Willard hypothesis (TWH) (Trivers and Willard 1973) implies that there are two possible cases in which fetuses are exposed to health shocks: scarring and selection.434343 Valente (2015) tests the TWH using the case of the civil conflict in Nepal and shows that fetal exposure to the war decreased the sex ratio at births.
Let $x\sim\mathcal{N}(\mu,\,\sigma^{2})$ be the initial health endowment of the fetus in utero, and assume that there is a survival threshold $x_{s}$ below which the fetus is culled before birth. [^44] Survival threshold is usually enough below the mean of the distribution as $x_{s}<\mu$. First, the scarring mechanism shifts the mean of the distribution ($\mu$) toward the left as $\mu^{*}<\mu$, such that $x^{*}\sim\mathcal{N}(\mu^{*},\,\sigma^{2})$. As the survival threshold $x_{s}$ is fixed, this shift implies that the number of culled fetuses must increase as $F^{*}(x_{s})>F(x_{s})$, where $F^{*}(\cdot)$ and $F(\cdot)$ indicate the cumulative distribution functions of the normal distributions with means $\mu^{*}$ and $\mu$, respectively. Similarly, the conditional expectation of the initial health endowment at birth satisfies the following condition:
$\footnotesize{\begin{split}E[x^{*}|x^{*}>x_{s}]<E[x|x>x_{s}].\end{split}}$ (4)
This means that the health endowment of surviving infants is more likely to decline due to fetal health shocks. [^45] See Online Appendix D.5 for the derivation of the conditional expectation of the truncated normal distribution.
Contrarily, the selection mechanism shifts the survival threshold ($x_{s}$) toward the right as $x_{s}^{*}>x_{s}$. As the mean of the distribution is fixed, this shift also increases the number of culled fetuses as $F(x_{s}^{*})>F(x_{s})$. Consequently, the conditional expectation of endowment at birth satisfies the following condition:
$\footnotesize{\begin{split}E[x|x>x_{s}]<E[x|x>x_{s}^{*}].\end{split}}$ (5)
This implies that the health endowment of surviving infants is more likely to be improved if the selection mechanism works.
The propositions can be summarized as follows. First, fetal health shocks may increase the risk of fetal death through both mechanisms. Second, the risk of infant mortality increases (decreases) in the scarring (selection) mechanism. Third, if both mechanisms work simultaneously, the risk of infant mortality remains unchanged. Considering these, the fetal death and infant mortality rates are used to test whether coal mines had adverse health effects and which mechanism worked in the case of coal mining operations.
Scarring due to the Occupational Hazards
Table 5: Testing the Occupational Hazards and Pollution
using the Early-life Mortality Data
Panel A: 1933 Vital Statistics
Analytical Sample
Full Treated Full Treated
(1) IMR (2) FDR (3) FDR (4) IMR (5) CMR (6) CMR
MineDeposit 36.74036.740**36.740 ∗ ∗ 21.19421.194**21.194 ∗ ∗ $1.460$
$(16.077)$ $(9.657)$ $(2.819)$
Female Miner 0.0380.038***0.038 ∗ ∗ ∗ 0.1040.104***0.104 ∗ ∗ ∗ $0.004$
$(0.012)$ $(0.031)$ $(0.004)$
Male Miner 0.003-0.003*- 0.003 ∗ 0.010-0.010**- 0.010 ∗ ∗ $-0.001$
$(0.002)$ $(0.004)$ $(0.001)$
River accessibility 2.122-2.122**- 2.122 ∗ ∗ $0.367$ $-2.993$ $-0.209$ 0.632-0.632***- 0.632 ∗ ∗ ∗ $0.603$
$(1.027)$ $(0.687)$ $(2.271)$ $(3.638)$ $(0.224)$ $(0.462)$
City and Town FEs Yes Yes Yes Yes Yes Yes
Railway accessibility Yes Yes Yes Yes Yes Yes
Port accessibility Yes Yes Yes Yes Yes Yes
Elevation Yes Yes Yes Yes Yes Yes
Observations 1,140 1,140 93 93 1,140 93
Estimator IV IV OLS OLS IV OLS
First-stage $F$-statistic $36.05$ $36.05$-- - --- - $36.05$-- -
Sample Mean of the DV $117.16$ $43.51$ $53.90$ $147.19$ $13.40$ $16.06$
Std. Dev. of the DV $42.66$ $27.41$ $23.54$ $43.95$ $9.58$ $5.54$
Panel B: 1938 Vital Statistics
Analytical Sample
Full Treated Full Treated
(1) IMR (2) FDR (3) FDR (4) IMR (5) CMR (6) CMR
MineDeposit $18.013$ $6.436$ $1.967$
$(12.025)$ $(7.952)$ $(2.462)$
Female Miner 0.0330.033**0.033 ∗ ∗ 0.0730.073***0.073 ∗ ∗ ∗ 0.0070.007*0.007 ∗
$(0.013)$ $(0.028)$ $(0.004)$
Male Miner $-0.000$ $-0.001$ $-0.000$
$(0.001)$ $(0.002)$ $(0.001)$
River accessibility 2.164-2.164**- 2.164 ∗ ∗ $0.429$ $0.227$ $-0.145$ $-0.332$ $-0.148$
$(1.013)$ $(0.575)$ $(1.996)$ $(3.485)$ $(0.213)$ $(0.701)$
City and Town FEs Yes Yes Yes Yes Yes Yes
Railway accessibility Yes Yes Yes Yes Yes Yes
Port accessibility Yes Yes Yes Yes Yes Yes
Elevation Yes Yes Yes Yes Yes Yes
Observations 1,364 1,364 115 115 1,364 115
Estimator IV IV OLS OLS IV OLS
First-stage $F$-statistic $44.95$ $44.95$-- - --- - $44.95$-- -
Sample Mean of the DV $114.33$ $37.30$ $50.16$ $135.91$ $16.81$ $20.70$
Std. Dev. of the DV $42.79$ $25.68$ $23.67$ $36.75$ $8.74$ $8.23$
***, **, and * represent statistical significance at the 1%, 5%, and 10% levels, respectively. Standard errors based on the heteroskedasticity-robust covariance matrix estimator are reported in parentheses.
Notes: Panels A and B show the results for the 1933 and 1938 Vital Statistics samples, respectively. The IMR, FDR, and CDR indicate the infant, fetal, and child mortality rates, respectively. Columns 1–2 show the results for the entire sample. Columns 3–4 show the results for the municipalities with coal mines (i.e., the municipalities included in the treatment group). Columns 5–6 show the results for the full sample and the municipalities with coal mines, respectively. Panel B of Table C.4 summarizes the summary statistics of the early-life mortality variables by treatment and control groups.
Panel A of Table 5 presents the results for the 1933 sample. Column 1 shows that the estimate for the IMR is positive and statistically significant, indicating that coal mines increased the infant mortality risk. In column 2, the same regression is run by replacing the IMR with FDR to assess the mechanism behind the higher risk of infant mortality in coal mining areas. The estimate is positive and statistically significant. This implies that greater risk was associated with reducing fetal health endowments by mining, supporting evidence on the scarring mechanism before birth.
Next, we test whether the occupational hazard led to this scarring due to coal mining. In column 3, the sample is limited to $93$ municipalities with coal mines to analyze the correlation between the FDR and the number of female miners and accessibility to rivers. If heavy manual work during pregnancy affects fetal health, FDRs should be positively correlated with the number of female miners. To control for the scale effect and provide a placebo test, the number of male workers was included in the same specification. The estimated coefficient for female miners is significantly positive, whereas that for male miners is negative. This result supports the evidence that occupational hazards for females (not males) alone increased the risk of death before birth. The estimate indicates that a one standard deviation increase in female miners ($=269.2$ miners) increases the number of fetal deaths by roughly $10$ ($0.038\times 269$) per 1000 births. This is economically meaningful because the mean difference in the FDR between the treatment and control groups was also approximately $9$ fetal deaths per 1000 births (Panel B of Online Appendix Table C.4). Column 4 shows the result for the IMR, showing that the estimated coefficient on the number of female miners is also statistically significantly positive. It further provides evidence that this reduction in fetal health endowment via occupational hazards for females is associated with higher infant mortality risks.
Panel B of Table 5 provides the results for the 1938 sample in the same panel and column layout. Columns 1–2 show positive but much smaller estimates, which are no longer statistically significant. The estimates listed in columns 3–4 are also positive, but smaller in magnitude. The estimate for the FDR shows that a one standard deviation increase in female miners ($=179.3$ miners) increased the number of fetal deaths by approximately $6$ ($0.033\times 179$) per 1000 births. This magnitude was roughly $60$% of that in 1933.
The decline in the estimated magnitude of these health outcomes between 1933 and 1938 is noteworthy. Although this could partly be associated with an improving trend in health-related risk-coping strategies such as increments in the number of medical doctors and installation of modern water supply (Online Appendix Figure D.5), the city and town fixed effects and accessibility variables can control for the heterogeneities in the local sanitary levels. Therefore, my results imply that the occupational hazards had increased the early-life mortality risk through the scarring mechanism in utero. However, the risks substantially decreased in the late 1930s because enforcing the revised labor regulations changed the labor patterns of females and reduced the number of female miners. Consequently, the institutional change may have attenuated the mortality selection mechanism before births throughout the 1930s.
Testing the Pollutions
Section 2 suggests that the coal sludge generated from the coal selection process could have been a potential pollution factor. If such wastewater increases the health risk of fetuses and infants, municipalities with mines closer to rivers should have significantly higher mortality rates because the wastewater had been discharged into the rivers (Chiba and Yamada 1964). However, the estimated coefficients of the river accessibility variable are close to zero and statistically insignificant in columns 3 and 4 in both panels. This result suggests that, while wastewater was occurring, it was not suffincient to harm the health conditions of mothers and infants, as anecdotes suggest (Section 2.3).
Moreover, air pollution is also unlikely to increase mortality risks. If air pollution increases early-life mortality, then the estimated coefficients on the male miner should be positive in columns 3 and 4. This is because larger mines should emit more coal smoke from the boilers of the steam winding machines. As shown, however, the estimates are very close to zero in all the cases. Column 5 of Panel A further tests the potential influence of air pollution using an alternative outcome, the mortality rate of children aged 1–4. If air pollution was a contributory factor, the child mortality rate would be greater in the coal mining area than in the surrounding regions. This is because younger children are more susceptible to pollutants and are more likely to play outside than infants. The estimate is, however, close to zero and is statistically insignificant. Column 6 of Panel A uses data on $93$ municipalities with coal mines and provides further evidence. The estimated coefficients on the number of female and male miners are close to zero, suggesting that child mortality did not depend on the scale of emissions. The results for child mortality in 1938 are materially similar to those in 1933 (columns 5–6 in Panel B). Overall, air pollution does not appear to be a plausible channel to explain the greater early-life mortality in the mining area. This is consistent with the historical fact that winding machines were electrified during the interwar period (Section 2.3).
5.5 Social Changes
Table 6: Testing the Mechanisms: Social Changes
Pre-regulation Post-regulation
Data: 1930 Census Data: 1935 Census
(1) Sex Ratio (2) HH Size (3) Sex Ratio (4) HH Size
MineDeposit 0.1150.115***0.115 ∗ ∗ ∗ 0.534-0.534***- 0.534 ∗ ∗ ∗ 0.0650.065***0.065 ∗ ∗ ∗ 0.486-0.486***- 0.486 ∗ ∗ ∗
$(0.034)$ $(0.141)$ $(0.023)$ $(0.140)$
City and Town FEs Yes Yes Yes Yes
Railway accessibility Yes Yes Yes Yes
Port accessibility Yes Yes Yes Yes
River accessibility Yes Yes Yes Yes
Elevation Yes Yes Yes Yes
Observations 1,140 1,140 1,364 1,364
Estimator IV IV IV IV
First-stage $F$-statistic $36.05$ $36.05$ $44.95$ $44.95$
Sample Mean of the DV $0.99$ $5.34$ $0.99$ $5.34$
Std. Dev. of the DV $0.08$ $0.47$ $0.07$ $0.57$
***, **, and * represent statistical significance at the 1%, 5%, and 10% levels, respectively. Standard errors based on the heteroskedasticity-robust covariance matrix estimator are reported in parentheses.
Notes: Columns 1–2 and 3–4 present the results for the 1930 and 1935 samples, respectively. Columns 1 and 3 show the results for sex ratio (male/female), whereas columns 2 and 4 show the results for average household size. Panel A of Online Appendix Table C.4 shows the summary statistics of these variables by treatment and control groups.
The fertility responses and mortality changes found in Sections 5.3 and 5.4 indicate that the natural change in the mining area was initially marginally positive given the low fertility and mortality differences with the surrounding area. After the enforcement of the revised regulation, the natural change may have become moderately positive because fertility reverted to a natural trend, and mortality slightly declined. Since short-run population dynamics is the sum of natural and social changes, social change shall be positive under the clear population growth in the mining area. Thus, estimating the number of immigrants is no longer relevant to understanding the mechanism behind the local population growth. Despite this, the extent to which the influx of new miners led to gender bias and influenced household size was analyzed in this subsection. This is useful to provide insight on how the fertility responses confirmed in Section 5.3 could be associated with the local demographic characteristics in the mining area.
Table 6 presents the results for the sex ratio and average household size. Columns 1–2 show the results for the 1930 sample, whereas Columns 3–4 show the results for the 1935 sample. The estimated coefficients in Columns 1 and 2 indicate that municipalities with coal mines experienced an increase in the sex ratio and a decline in the average household size. The estimated magnitudes are economically meaningful. Both estimates ($0.12$ and $0.54$) account for more than one standard deviation in the sex ratio and the average household size, respectively. This confirms that the influx of miners, many of whom were single males, led to a disproportionate sex ratio and smaller household sizes. [^46] This tendency in the mining area is in line with the phenomenon observed in the urban areas in prewar Japan. The urban areas had a relatively higher sex ratio than rural localities due to the internal migration of male workers (Ito 1990, p.243–247). Columns 3–4, however, show much smaller estimates for the 1935 sample. The estimates for the sex ratio and average household size are $0.07$ and $0.49$. While both estimates are still statistically significant, the large declines in magnitude are relevant to the reverting fertility after the regulation (Section 5.3). The increased fertility reverses the overall sex ratio because the sex ratio at birth is generally balanced at that time. It also increases the average household size of existing households, attenuating the impacts of single immigrants.
To summarize, the foregoing results provide evidence suggesting that the miners had begun to form their families in the mid-1930s in response to the labor regulations targeted at female miners. The total number of miners had been relatively stable between 1930 and mid-1930s (Table 1), suggesting that the social changes in the mining areas were stable. Consequently, a possible explanation for the greater marginal effect on the local population growth in the mid-1930s in the main result (Table 2) is that an increase in the natural growth term by the local family formation and declines in the mortality risk through the institutional change in the early 1930s.
5.6 Robustness
The robustness of the results is ensured in the following ways. First, the cluster-robust variance-covariance matrix estimator was used in Online Appendix D.3. This estimator provides a systematically larger estimate for the standard error by allowing for within-cluster dependencies. Despite this, the results are materially similar, meaning that the baseline estimates are robust to the potential influences of the local-scale spatial correlations. Second, the regional heterogeneity in the endowment of male miners is taken into account. The coal mines in Hokkaidō (northernmost island) were less likely to use female miners in the pits, given the nature of the labor supply of male miners. The results are materially similar if we include an indicator variable for Hokkaidō prefecture in the regressions (Online Appendix D.7). Third, an alternative identification strategy is employed. As explained in the data section, the nature of the assignments makes it difficult to employ the estimation strategy utilizing within variations of each unit. Nevertheless, the case study is provided using the DID approach for the subsample of the Tōhoku region (northeastern region). The results are reasonable to the main findings (Online Appendix D.8). Finally, the validity of the results on coal mines is assessed by comparing the results for heavy metal mines. Since coal mining requires many more workers than heavy metal mining because coal is a bulky mineral, the estimated magnitudes for the heavy metal mines shall be much smaller. It is confirmed that the results are reasonable to such expectation (Online Appendix D.9).
6 Conclusion
The innovation of labor-saving technologies occurred with the renewal of traditional extraction methodology in the coal mining sector during interwar Japan. These technology shocks induced the revision of labor regulations, leading to institutional change concerning the female mines. In this study, the impact of coal mines on the regional economy during historical industrialization in interwar Japan is examined to investigate how the institutional change influenced the effect.
In order to avoid selection bias, a set of official census-based statistics was used. The exogenous variation from the geological stratum is exploited to structurally identify the impacts of the coal mines on the regional development. Several robustness checks confirm that the main results are not sensitive to the potential confounding factors, influential observations, and the use of alternative variation in the estimation.
It is found that the coal mines increased the local population, and the estimated magnitude is greater in the sample after the full enforcement of the revised labor regulations. Evidence from the analyses using the register-based vital statistics indicates that the regulations had indeed forced female miners to exit the labor market and form families. Before the full enforcement of the regulation, the occupational hazards for female miners substantially increased the early-life mortality via the biological mortality selection mechanism in utero. The institutional change, however, further led to improvements in early-life mortality.
The literature on institutional interventions in the labor market has provided evidence that labor regulations can impede regional developments as they can reduce the local labor supply and increase unemployment (Botero et al. 2004). The revised labor regulations reduced the female labor supply and led to the gender-biased structural shift during interwar Japan. The findings, however, show that technology-induced labor regulations accelerated the local population growth through the family formation channel. In this process, the regulations played a role in mitigating the occupational hazards for female miners and improving the early-life health status in the local economy. The results demonstrate the role of institutional change from gender-specific labor regulations in understanding the mechanisms behind regional developments.
Appendices
Appendix Appendix A Background Appendix
A.1 Influence of Great Depression in Japan
The impact of the Great Depression was not substantial in Japan compared to other countries (Miyamoto 2008, pp. 56–57; Abe 2008, p. 106; Blumenthal 1981, pp. 46). Blumenthal (1981, p.43) shows that the index of industrial production in Japan did not decrease dramatically between 1929 (=100) and 1930 (=97). In contrast, the US and the UK suffered much more significant reductions (81 in the US and 92 in the UK) during the same period. Blumenthal (1981, p. 45) also reports that the unemployment rates between 1930 and 1935 in the US, the UK, and Japan were approximately 27%, 15%, and 6%, respectively. Consequently, Japan’s GNP growth rate (from 1929 to 1930) was positive at $1.1$%, whereas it was $-7.7$ and $-0.1$, respectively, in the US and UK. This evidence supports the validity of using the systematic data from the census and vital statistics of the 1930s. In the mining sector, while several mines in Chikuhō coalfield were documented to have ceased operations during the depression, these were temporary (Compilation Committee of the Chronology of Chikuhō Coal Mining Industry 1973). Figure 1 confirms that the reduction in coal production was marginal in the case of Japan. Overall, the depression mainly hit prices rather than production. While the price declines were relatively large in textiles, metals, and construction materials, the shocks influenced food and fuel prices less (Nakamura and Odaka 1989, p. 53).
A.2 Female Miners in Coal Mine
By the end of the 19th century, many mines in the Chikuhō coalfields had finished mechanizing the coal haulage processes. However, this did not mean that the miner’s workload related to hauling operations was reduced. Instead, this mechanization in the main shafts had increased the workload of miners extracting and transporting coal around the coalface (Sumiya 1968, pp. 310–312).
Figure A.1a shows a photograph of miners at the mine face (kiriha) in the 1910s. A male skilled miner (saki-yama) extracted coal, whereas a female miner (ato-yama) brought the coal to the surface using a bamboo basket called sura. Figure A.1b shows that ato-yama tied a rope around her body to prevent the heavy sura from slipping off in the steep pits. The environment in the pit was generally poor. Moreover, accidents caused by falling rocks occurred frequently. [^47] Available statistics indicate that the average number of accidents per year that occurred in the pits between 1917 and 1926, causing injuries and deaths of miners, was 142,595, of which 61,112 were from falling rocks. (Japan Mining Association, 1928, p. table.6). Accidents caused by explosions, poisoning, and suffocation were fewer than those caused by falling rocks, but they still occurred 311 times per year on average between 1917 and 1926 (Japan Mining Association 1928, p. table.6)..
(a) Saki-yama (left) and Ato-yama (right)
(b) Ato-yama using Sura
Figure A.1: Male (saki-yama) and Female Miners (ato-yama)
Notes: Figure A.1a shows the skilled male miner (saki-yama) extracting coal and the female miner (ato-yama) supporting her husband at the face. Figure A.1b illustrates an ato-yama bringing coal using a bamboo basket (sura) at the gangway through the steep pits. Both figures were taken in the early Taishō period.
Source: Fumoto (1961, p. 2).
The introduction of new mining technologies, such as coal cutters, was gradually adopted in the 1920s. The introduction of these technologies was accompanied by a transition from the traditional room-and-pillar mining method to the longwall method, which is more suitable for mechanical coal extraction. [^48] Compared to the room-and-pillar method, the coal extraction space of the longwall method is larger, allowing for the introduction of larger machinery (Nishinarita 1985, p. 91). The number of miners, both male and female, declined during this transition. However, this did not entirely change the labor patterns of female miners. Approximately one in five miners was a woman in 1930 (Table 1), and they still worked inside the pits as ato-yama. [^49] See also Tanaka (1984, chapter. 4) for a comprehensive summary of the relationships between mechanization process and changes in labor patterns in the pits. In short, introducing coal conveyors was the primary driving force for replacing the in-pit works of the female mines. This provided a strong incentive for mining companies to exclude female miners from the pits (Tanaka and Ogino 1977, p. 79). Finally, the Revised Miners’ Labor Assistance Regulations of September 1928 prohibited some underground and late-night work. The number of female miners in the pits (ato-yama) dropped from 1930 to 1933, when the Revised Miner’s Labor Assistance Regulations were fully enacted in September (Table 1). Importantly, however, the revised regulations did not eliminate female miners altogether. Female miners were allowed to work in the pits of the thin layer and were primarily responsible for coal selection and other out-of-pit labor. More than 15,000 women still worked in the mines in the late 1930s, representing approximately 10% of all the miners measured.
A.3 Mechanization and Female Miners
The history illustrated in Online Appendix A.2 suggests that female miners were more likely to be replaced via mechanization. If this anecdote is accurate, the mines deploying more machines were less likely to rely on the female miners. To assess this point, I examine the relationship between the degree of mechanization and the use of female miners in coal mines. The systematic data capturing mechanization in coal mines is unavailable. Fortunately, Fukuoka Prefectural Statistics document the horsepower and the number of miners in the coal mines in Fukuoka, including the Chikuhō coalfield. Figure A.2 illustrates the raw relationship between the share of female miners and log-transformed horsepower in the 82 coal mines in the prefecture in 1930. Overall, this confirms the negative correlation between both variables. The estimate from the Tobit model dealing with several censored observations in Figure A.2 is $-4.197$, which is statistically significant at the $1$% level. This suggests that a one standard deviation increase in the log-transformed horsepower ($=2.0$) decreases the use of female mines by approximately $8.4$%, which accounts for more than half of the standard deviation of the share of female miners ($=15.6$). Thus, this result supports the idea that mechanization via the installation of new technologies has worked to replace female miners in the coal mines.
Figure A.2: Relationship between share of female miners (%) and log-horsepower (hp) in coal mines in Fukuoka
Note: This figure illustrates the raw relationship between the share of female miners and log-transformed horsepower in 82 coal mines in Fukuoka prefecture in 1930. The share of female miners is the number of female miners per 100 total miners. The mean and standard deviation of the share of female miners are $20.1$ and $15.6$, respectively. The horsepower is the total horsepower of the machines using any of steam, electrical, and water power. The mean and standard deviation of the log-transformed horsepower are $7.1$ and $2.0$, respectively. The red line represents the fitted values for a linear regression model. Although the Fukuoka Prefectural Statistics originally contained 90 coal mines, 8 mines were dropped because there were missing values in the horsepower statistics.
Source: Created by the author. Data are obtained from the official report published by Fukuoka Prefecture (1932).
A.4 Shares of Coal and GSC Mines
(a) Coal and oil mines
(b) Metal mines
Figure A.3: The Number of Coal, Oil, GSC, and the Other Metal Mines
in 1931 and 1936
Notes: This figure shows the number of coal and oil (Figure A.3a) and metal mines (Figure A.3b) documented in the AHCs of 1931 and 1936. GSC indicates gold, silver, and copper mines.
Source: Created by the author. Data on mines are from the Association for Harmonious Cooperation (1932, 1937).
Figure A.3a shows the number of coal and oil mines documented in the AHCs in 1931 and 1936. Similarly, Figure A.3b shows the numbers of GSC and other metal mines. As shown, approximately 90 (85)of mines extracting fuels (metals) were coal (GSC) mines. Figure A.3 shows the number of mines. Since there were some cases where a few mines coexist in a municipality, the number of municipalities with mines reported in the main text is smaller than the number of mines.
Appendix Appendix B Conceptual Framework Appendix
The regulations of female miners decreased the relative wage rate of female miners (Section 2.2). The purpose of this section is to provide a conceptual framework for how their fertility changes in response to the female relative wages. I consider a simplified case of a collective model proposed by Siegel (2017) that can illustrate the optimal fertility based on the couple household’s optimization.
Setting
Males and females derive utilities from consumption ($c$) and having children ($b$). For a spouse $g\in\{m,f\}$, additively separable utility function is specified as follows:
$\footnotesize{\begin{split}u_{g}(c_{g},b)=\log c_{g}+\tau\log b.\end{split}}$ (6)
Home production required to raise children can be expressed as:
$\footnotesize{\begin{split}y(b)=\pi b,\end{split}}$ (7)
where $\pi>0$. The home production is accomplished using labor-saving technologies available in the markets ($s$) and parents’ total home labor input ($\Omega$):
$\footnotesize{\begin{split}y(b)=s^{\gamma}\Omega^{1-\gamma},\end{split}}$ (8)
where $10\leq\gamma<$ 1 represents the availability of childcare services in the economy. Let $\kappa_{g}$ be the productivity of houseworks including childcare, and $h_{g}\geq 0$ be the home labor input of each spouse. The parents’ total home labor input is then specified in the constant elasticity of the substitution function as:
$\footnotesize{\begin{split}\Omega=\left(\kappa_{m}h_{m}^{1-\rho}+\kappa_{f}h_{f}^{1-\rho}\right)^{\frac{1}{1-\rho}},\end{split}}$ (9)
where $\rho>0$ is the inverse of the elasticity of substitution. A spouse earns wage $w_{g}$ for labor input $n_{g}\geq 0$. The labor income is used for their consumption and for purchasing home labor-saving services.
Optimization
The representative household maximizes the sum of utilities:
$\footnotesize{\begin{split}\theta u_{m}(c_{m},b)+(1-\theta)u_{f}(c_{f},b)\end{split}}$ (10)
subject to the budget and time constraints:
$\footnotesize{\begin{split}c_{m}+c_{f}+s=w_{m}n_{m}+w_{f}n_{f},\\ n_{g}+h_{g}=1~{}\text{for}~{}g\in\{m,f\}.\end{split}}$ (11)
The optimal home labor inputs and fertility are given by:
$\footnotesize{\begin{split}h_{m}=\left(\frac{\kappa_{m}}{w_{m}}\right)^{\frac{1}{\rho}}\left(\frac{1-\gamma}{\gamma}\right)^{\gamma}\{\kappa_{m}^{\frac{1}{\rho}}w_{m}^{\frac{1}{\rho-1}}+\kappa_{f}^{\frac{1}{\rho}}w_{f}^{\frac{1}{\rho-1}}\}^{\frac{\gamma\rho-1}{1-\rho}}y_{b},\end{split}}$ (12)
$\footnotesize{\begin{split}h_{f}=\left(\frac{\kappa_{f}}{w_{f}}\right)^{\frac{1}{\rho}}\left(\frac{1-\gamma}{\gamma}\right)^{\gamma}\{\kappa_{m}^{\frac{1}{\rho}}w_{m}^{\frac{1}{\rho-1}}+\kappa_{f}^{\frac{1}{\rho}}w_{f}^{\frac{1}{\rho-1}}\}^{\frac{\gamma\rho-1}{1-\rho}}y_{b},\end{split}}$ (13)
$\footnotesize{\begin{split}b=\frac{\tau}{\pi}\left(\frac{c_{m}}{\theta}\right)\gamma^{\gamma}\left(1-\gamma\right)^{1-\gamma}\{\kappa_{m}^{\frac{1}{\rho}}w_{m}^{\frac{1}{\rho-1}}+\kappa_{f}^{\frac{1}{\rho}}w_{f}^{\frac{1}{\rho-1}}\}^{\frac{\rho(1-\gamma)}{1-\rho}}.\end{split}}$ (14)
Equations 12 and 13 characterize the optimal ratio of home labor inputs as $\frac{h_{m}}{h_{f}}=\left(\frac{\kappa_{m}}{\kappa_{f}}\frac{w_{f}}{w_{m}}\right)^{\frac{1}{\rho}}$. Substitute the optimal consumption and optimal time-allocation choice into budget constraints to yield:
$\footnotesize{\begin{split}\frac{c_{m}}{\theta}=(w_{m}+w_{f})-\left(\frac{1}{\gamma}\right)^{\gamma}\left(\frac{1}{1-\gamma}\right)^{1-\gamma}\{\kappa_{m}^{\frac{1}{\rho}}w_{m}^{\frac{1}{\rho-1}}+\kappa_{f}^{\frac{1}{\rho}}w_{f}^{\frac{1}{\rho-1}}\}^{\frac{\rho(1-\gamma)}{\rho-1}}y_{b}.\end{split}}$ (15)
This implies that equation 14 can be written as a closed form:
$\footnotesize{\begin{split}b=\frac{\tau}{\pi(1+\pi)}\gamma^{\gamma}\left(1-\gamma\right)^{1-\gamma}\left(w_{m}+w_{f}\right)\{\kappa_{m}^{\frac{1}{\rho}}w_{m}^{\frac{1}{\rho-1}}+\kappa_{f}^{\frac{1}{\rho}}w_{f}^{\frac{1}{\rho-1}}\}^{\frac{\rho(1-\gamma)}{1-\rho}}.\end{split}}$ (16)
This is independent of the bargaining weight ($\theta$). Finally, equation 16 suggests that $\frac{\partial b}{\partial w_{f}}$ becomes negative under following condition:
$\footnotesize{\begin{split}\gamma\kappa_{f}^{\frac{1}{\rho}}w_{f}^{\frac{\rho-1}{\rho}}+w_{m}\left[\left(\frac{\kappa_{m}}{w_{m}}\right)^{\frac{1}{\rho}}-(1-\gamma)\left(\frac{\kappa_{f}}{w_{f}}\right)^{\frac{1}{\rho}}\right]<0.\end{split}}$ (17)
Prediction
Since prewar Japan was a paternalistic society under the Old Civil Code, the most plausible setting for housework productivity at that time would be $\kappa_{m}=0$. Under this setting, inequality 17 indicates that $\frac{\partial b}{\partial w_{f}}$ is negative under the following condition:
$\footnotesize{\begin{split}\frac{w_{f}}{w_{m}}<\frac{1-\gamma}{\gamma}.\end{split}}$ (18)
This suggests that female wages negatively correlate with fertility when adequate childcare services are unavailable in local markets.
Table B.1: Nursary at Yamauchi Coal Mine in 1915
(a) Ave. years of (b) Ave. years of (c) Ave. days using (d) Ave. days (e) Utilization
service of parent use of nursery ($\bar{p}_{c}$) nursary ($\bar{n}_{c}$) of operation rate (%)
2.4 1.3 85.2 320 20.0
(1.3) (1.0) (57.5)
Notes: Columns (a) to (c) are from the survey report on $88$ children using the nursery at Yamauchi Coal Mine in Chikuhō coalfield in 1915. The figures in parentheses are the median values. Column (d) shows the average number of operating days of coal mines in 1924, surveyed in the mining census. The average number of operating days is calculated as 365 minus the average number of days closed ($45.5$). Column (e) shows the utilization rate, which is defined as: $100\left(\sum_{c=1}^{88}n_{c}\right)/\left(\sum_{c=1}^{88}p_{c}\times(365-45.5)\right)$.
I excluded several individuals whose statistics were unavailable or unreliable. Three individuals whose length of service was longer than the number of years they had used the nursery are excluded. Six individuals had more days in care than days in operation. These individuals are also excluded because of suspected listing and transcription errors. Finally, I excluded three individuals for whom statistics on the number of days of childcare were not available. Consequently, this table uses statistics on the 88 individuals among $100$ individuals originally listed. Sources: Noyori 2010, pp. 80–82. The average number of days closed is obtained from the Statistics Bureau of the Cabinet 1926, p. 18.
To what extent were childcare services available to female miners at that time? As described in Section 2, female miners could use the nurseries attached to some coal mines. Fortunately, Noyori (2010) provides a comprehensive study on the nurseries in coal mines using a unique archive of the nursery survey in a few mines in Chikuhō coalfield. According to the nursery’s regulations, employees can use the nursery during working hours (Noyori 2010, p. 78). However, evidence shows that children over 5 years old were unlikely to use the nurseries in mines (Noyori 2010, pp. 76–77). Moreover, the number of days using the nursery was quite small. Table B.1 summarizes the utilization of the nursery based on the survey report on 88 children at the Yamauchi Coal Mine in 1915. Column (a) shows that mothers’ average working years are 2.4, whereas their average years of nursery use are 1.3. Column (c) indicates that the average number of annual days used was roughly 85 days (median is $57.5$ days), whereas the average number of annual operating days available from the mining census is 320 (column d). The years spent in daycare ($p_{c}$) can be considered the number of years the mother worked after the child was born. Thus, one can calculate the utilization rate for each child ($c$) as follows:
$\footnotesize{\begin{split}\frac{\sum_{c=1}^{88}n_{c}}{\sum_{c=1}^{88}p_{c}\times 320}\times 100,\end{split}}$ (19)
where $n_{c}$ is the number of days using nursery. Column (e) indicates that this rate is estimated to be $20$%. Note that several mothers used the nursery for more than 200 days, meaning that the nursery must have been opened throughout the operating days. This rejects the possibility that the nursery could not be opened due to a lack of staff. Despite this, it is indeed possible that miners might not work fully during the operating days. Even if one assumes that miners worked only one-third of the operating days (i.e., $107$ days), the utilization rate is still $60$%.
In addition, childcare requires fees, which were usually deducted from miners’ paychecks (Noyori 2010, p. 109). This implies that some miners who could not use the nursery may use alternative childcare institutions. Another possible informal institution may be hiring babysitting labor (komori). Noyori (2010, pp. 47–48) shows that these babysitters were usually children from other households. The young babysitter was expected to carry the infant on her back up to the mine entrance. The mother would then climb up and down multiple times like a “mouse” from the pits to the entrance to feed her baby. This anecdotal evidence suggests that the quality of babysitting was not high enough to compensate for childrearing.
Overall, while the nursery was a useful institution for female miners, it only replaced a small part of the childcare work of female miners. Although babysitting was available, it offered an incomplete service, covering only a limited portion of mothers’ childcare needs. This meant that the marketization parameter $\gamma$ would be considerably low in inequality 18. Therefore, optimal fertility should be negatively correlated with the relative wage ($w_{f}/w_{m}$). This predicts the relationship that the fertility rate would be lower before the female labor regulation but could be higher after the regulation. This is because the relative wage declined substantially after the prohibition of female miners’ work in the pits (Section 2.2).
Appendix Appendix C Data Appendix
C.1 Mine Deposit
Data on the location and type of mines are obtained from the official reports published by the Association for Harmonious Cooperation (1932, 1937). They document all mines with more than $50$ miners in October 1931 and October 1936. The documents list three types of mines: metal, coal, and oil. [^50] The metal mine includes minerals such as gold, silver, copper, blister copper, pig iron, steel, iron sulfide, lead, zinc, bismuth, arsenious acid, tin, mercury, chromium, manganese, and sulfide minerals. A smelting mine is also included in the metal mine category. Coal mines include a few mines mining lignite. Oil mines include oil, crude oil, gas, volatile oil, kerosene, light oil, machine oil, and heavy oil. In these documents, the mine’s location is indicated by the name of the municipality. Hence, the name of the municipality was used while geocoding to match the mine dataset with the datasets of the other variables. The shapefile used for geocoding was obtained from the official database of the Ministry of Land, Infrastructure, Transport and Tourism (available at: https://nlftp.mlit.go.jp/ksj/gml/datalist/KsjTmplt-N03-v2˙2.html (accessed on 23rd February 2021). Several municipalities were consolidated between the years 1930 and 1938. Thus, the municipality boundaries defined in 1938 were used and aggregate statistics for consolidated municipalities. A minimal number of municipalities were divided; these were excluded from the original dataset. To check all consolidations and divisions of municipalities, Zenkoku shichōsonmei hensen sōran (Municipal Autonomy Research Association 2006), was used because it is the most comprehensive book on the history of municipalities’ names.
The initial dataset included both coal and heavy metal (gold, silver, and copper) mines. [^51] This is because the most representative heavy metals at that time in the Japanese mining sector were gold, silver, and copper (Online Appendix D.9). Therefore, to prepare the analytical samples, the duplications were excluded to better identify the impact of coal mines. Panel A of Table C.1 summarizes duplications in the coal mine sample. Columns (1) and (2) show the figures based on the 1931 (1936) AHC, which is used to match with the 1930 (1935) Population Census and 1933 (1938) Vital Statistics datasets. Initially, there are $11,151$ municipalities in columns (1) and (2); note that the municipality boundaries are defined based on the 1938 boundaries. The number of municipalities within 30km of a coal mine is $1,147$ and $1,386$ in columns (1) and (2), respectively. The number of duplications in the treatment groups based on the $5$ km radius definition is 00 and $2$ in columns (1) and (2), respectively. Similarly, the number of duplications in the control groups based on the $5$ km radius definition is $7$ and $20$ in columns (1) and (2), respectively. Following the trimming step shown in Figure C.1a, the analytical coal mine samples are $1,140$ ($=1,147-7$) and 1,364 ($=1,386-22$), respectively. Panel B of Table C.1 summarizes the duplications in the GSC mine sample. Following the same procedure as in Figure C.1b, the GSC mine samples used in the robustness analysis are $2,498$ ($=2,510-22$) and $3,293$ ($=3,337-44$), respectively.
The red (blue) circles in Figure C.2 indicate the coal (GSC) mine locations from the AHC. Figure 2 illustrates the spatial distribution of the coal mine sample by year and group. Figure 3 shows the distribution in Chikuhō coalfield in Kyūshū in detail.
Table C.1: Duplications between Coal and GSC Mines
(1)1930 or 1933 (2)1935 or 1938
Panel A: Coal mine sample
Number of total municipalities 11,151 11,151
Number of municipalities within 30km from a coal mine 1,147 1,386
Duplications in treatment group
      Number of municipalities with a coal mine 93 115
          of which has a GCS mine 0 1
      Number of municipalities within 5km from a coal mine 167 179
          of which has a GCS mine within 5km 0 2
Duplications in control group
      Number of municipalities with no coal mines 1,054 1,335
          of which has a GCS mine 3 12
      Number of municipalities with no coal mines within 5km 980 1,271
          of which has a GCS mine within 5km 7 20
Panel B: GSC mine sample
Number of total municipalities 11,151 11,151
Number of municipalities within 30km from a coal mine 2,510 3,337
Duplications in treatment group
      Number of municipalities with a GSC mine 67 111
          of which has a coal mine 0 1
      Number of municipalities within 5km from a coal mine 125 174
          of which has a coal mine within 5km 0 2
Duplications in control group
      Number of municipalities with no GSC mines 2,437 3,196
          of which has a coal mine 6 29
      Number of municipalities with no GCS mines within 5km 2,373 3,119
          of which has a coal mine within 5km 12 42
Source: See Online Appendix C.1.
(a) Coal
(b) GSC
Figure C.1: Construction of Analytical Samples
Notes: These figures demonstrare the procedure to trim the municipalities into the analytical samples. Figures C.1a and C.1b show the trimming steps for the coal and GSC (gold, silver, and copper) mines, respectively.
Sources: Created by the author.
(a) 1931
(b) 1936
Figure C.2: Location of Coal and GSC mines in 1931 and 1936
Notes: The red and blue circles indicate the locations of coal and GSC (gold, silver, and copper) mines, respectively.
Source: Created by the author. Data on the location of mines are from the Association for Harmonious Cooperation (1932, 1937).
C.2 Geological Stratum
[…]
Tn and Tp are sedimentary rocks created during the Cenozoic era (shinseidai). Tp was created during a much earlier period, called the Paleogene period (ko daisanki), compared to Tn, which was created during the Neogene period (shin daisanki). Evidence indicates that Tp includes coal layers created in the carboniferous ages (Nagao 1975; Tanaka 1984). Figure C.4 shows a geological columnar section of the Iizuka coal mine, a representative coal mine in the Chikuhō coalfield. This figure provides evidence that many coal layers are included in the layers created in the Paleogene period.
Figure C.3: Spatial Distribution of Mines and Geological Strata
Note: The white circles indicate the location of the mines. The strata are illustrated through colors.
Sources: Created by the author. Data on the location of mines are obtained from the Association for Harmonious Cooperation (1932, 1937). Data on the geological stratum are obtained from the Ministry of Land, Infrastructure, Transport and Tourism.
Table C.2: Defining Instrumental Variables for the Coal Mine Subsample
Mining points in 1931 Mining points in 1936
Stratum Coeff. Std. Err. $R$-squared Coeff. Std. Err. $R$-squared
      Sedimentary rock
          M $0.060$ $(0.073)$ $0.001$ $0.025$ $(0.051)$ $0.000$
          P $0.036$ $(0.122)$ $0.000$ $-0.034$ $(0.055)$ $0.000$
          Pls $0.020$ $(0.167)$ $0.000$ $0.037$ $(0.167)$ $0.000$
          Tn 0.1430.143***0.143 ∗ ∗ ∗ $(0.040)$ $0.017$ 0.1930.193***0.193 ∗ ∗ ∗ $(0.042)$ $0.028$
[…]
          ds 0.147-0.147***- 0.147 ∗ ∗ ∗ $(0.011)$ $0.001$ 0.131-0.131***- 0.131 ∗ ∗ ∗ $(0.009)$ $0.001$
          lm $0.020$ $(0.167)$ $0.000$ $-0.072$ $(0.060)$ $0.001$
          tr 0.106-0.106***- 0.106 ∗ ∗ ∗ $(0.020)$ $0.015$ 0.075-0.075***- 0.075 ∗ ∗ ∗ $(0.019)$ $0.008$
          wf 0.123-0.123***- 0.123 ∗ ∗ ∗ $(0.030)$ $0.004$ 0.109-0.109***- 0.109 ∗ ∗ ∗ $(0.026)$ $0.003$
      Metamorphic rock
          Gn 0.147-0.147***- 0.147 ∗ ∗ ∗ $(0.011)$ $0.000$ 0.132-0.132***- 0.132 ∗ ∗ ∗ $(0.009)$ $0.002$
          Sch $-0.011$ $(0.046)$ $0.000$ $-0.007$ $(0.042)$ $0.000$
      Body of water
          w 0.147-0.147***- 0.147 ∗ ∗ ∗ $(0.011)$ $0.001$ 0.131-0.131***- 0.131 ∗ ∗ ∗ $(0.009)$ $0.001$
      Igneous rock
          An 0.108-0.108***- 0.108 ∗ ∗ ∗ $(0.025)$ $0.007$ 0.107-0.107***- 0.107 ∗ ∗ ∗ $(0.020)$ $0.007$
          Bs $-0.031$ $(0.051)$ $0.000$ $-0.027$ $(0.046)$ $0.000$
          Gd $0.133$ $(0.051)$ $0.002$ $0.122$ $(0.112)$ $0.002$
          Gr 0.059-0.059**- 0.059 ∗ ∗ $(0.025)$ $0.004$ 0.058-0.058***- 0.058 ∗ ∗ ∗ $(0.021)$ $0.004$
          Lp 0.150-0.150***- 0.150 ∗ ∗ ∗ $(0.011)$ $0.004$ 0.134-0.134***- 0.134 ∗ ∗ ∗ $(0.009)$ $0.005$
          Sp $0.054$ $(0.200)$ $0.000$ $0.013$ $(0.143)$ $0.000$
***, **, and * represent statistical significance at the 1%, 5%, and 10% levels, respectively. Standard errors based on the heteroskedasticity-robust covariance matrix estimator (Horn et al. 1975) are reported in parentheses.
Notes: This table summarizes the results of the linear probability models for the $19$ types of strata. The number of observations for 1931 and 1936 are $1,140$ and $1,364$, respectively.
Figure C.4: Geological Columnar Section of the Iizuka Coal Mine
Note: This geological columnar section shows the strata around the Iizuka coal mine in Chikuhō coalfield in Kyūshū region. The first to third columns include the names of layers created in the Paleogene period. The fourth to seventh columns examine the length of each layer, columnar, and legend. The black layers shown in the columnar of the sixth column are the coal layers: more than ten coal layers are included in the layers created in the Paleogene period.
Source: Fumoto (1961, p. Chishitsu chūjyō zu).
Table C.3: Summary Statistics: Mine and Instrumental Variables
Year Coal mines GSC mines
Mean Std. Dev. Obs. Mean Std. Dev. Obs.
Mine Deposit 1931 $0.146$ $0.354$ $1140$ $0.050$ $0.218$ $2498$
1936 $0.130$ $0.336$ $1364$ $0.053$ $0.224$ $3293$
Stratum (1931) $0.094$ $0.292$ $1140$-- - --- - --- -
(1936) $0.086$ $0.280$ $1364$-- - --- - --- -
Female Miners 1931 $207.51$ $269.16$ $93$ $51.91$ $71.13$ $57$
1936 $145.65$ $179.77$ $115$ $40.73$ $59.18$ $111$
Male Miners 1931 $1359.25$ $1710.20$ $93$ $511.86$ $784.73$ $57$
1936 $1512.84$ $2040.87$ $115$ $451.01$ $680.58$ $111$
Notes: This table shows the summary statistics for the indicator variable for the treatment group, geological stratum, and number of miners. The treatment (control) group is 0–5 (5-30) km from the centroid of a municipality that has a mine. The geological stratum is a time-constant variable: “Year” in parenthesis indicates the matched year. GSC mines indicate gold, silver, and copper mines.
Sources: Data on the location of mines are from the Association for Harmonious Cooperation (1932, 1937).
C.3 Dependent Variables
Table C.4: Summary Statistics: Outcome Variables
Panel A: Year Overall Treated Controlled
Demographics Mean Std. Dev. Obs. Mean Std. Dev. Obs. Mean Std. Dev. Obs.
Population 1930$\ddagger$ $6305.06$ $14829.74$ $1140$ $13091.93$ $25577.78$ $167$ $5140.20$ $11692.06$ $973$
1935$\ddagger$ $6202.12$ $15598.06$ $1364$ $12986.87$ $28001.89$ $177$ $5190.41$ $12464.17$ $1187$
Sex ratio (male/female) 1930$\ddagger$ $0.99$ $0.08$ $1140$ $1.03$ $0.10$ $167$ $0.98$ $0.07$ $973$
1935$\ddagger$ $0.99$ $0.07$ $1364$ $1.03$ $0.08$ $177$ $0.98$ $0.07$ $1187$
Household size 1930$\ddagger$ $5.34$ $0.47$ $1140$ $5.18$ $0.41$ $167$ $5.37$ $0.47$ $973$
1935$\ddagger$ $5.34$ $0.57$ $1364$ $5.20$ $0.44$ $178$ $5.36$ $0.59$ $1187$
Crude marriage rate 1930$\ddagger$ $8.59$ $2.28$ $1140$ $7.43$ $2.19$ $167$ $8.79$ $2.24$ $973$
1935$\ddagger$ $9.04$ $2.41$ $1364$ $7.98$ $2.05$ $177$ $9.20$ $2.42$ $1187$
Crude birth rate 1930$\ddagger$ $34.06$ $5.52$ $1140$ $31.75$ $5.62$ $167$ $34.46$ $5.40$ $973$
1935$\ddagger$ $34.22$ $5.47$ $1364$ $32.37$ $4.96$ $177$ $34.50$ $5.49$ $1187$
Marital fertility rate 1930$\ddagger$ $174.14$ $29.69$ $1140$ $161.91$ $30.58$ $167$ $176.24$ $29.04$ $973$
1935$\ddagger$ $177.50$ $29.75$ $1364$ $169.40$ $27.87$ $177$ $178.70$ $29.84$ $1187$
Crude death rate (male) 1930 $20.06$ $4.48$ $1140$ $20.04$ $4.62$ $167$ $20.06$ $4.45$ $973$
1935 $18.36$ $4.12$ $1364$ $177$ $17.90$ $177$ $18.43$ $4.14$ $1187$
Crude death rate (female) 1930 $18.59$ $4.33$ $1140$ $18.20$ $3.88$ $167$ $18.66$ $4.40$ $973$
1935$\dagger$ $16.97$ $4.21$ $1364$ $16.30$ $3.89$ $177$ $17.08$ $4.24$ $1187$
Panel B: Year Overall Treated Controlled
Early-life health Mean Std. Dev. Obs. Mean Std. Dev. Obs. Mean Std. Dev. Obs.
Infant mortality rate 1933$\ddagger$ $117.15$ $42.66$ $1140$ $139.40$ $44.56$ $167$ $113.34$ $41.15$ $973$
1938$\ddagger$ $114.33$ $42.79$ $1364$ $131.97$ $36.79$ $177$ $111.70$ $43.02$ $1187$
Fetal death rate 1933$\ddagger$ $43.51$ $27.41$ $1140$ $51.57$ $26.25$ $167$ $42.13$ $27.38$ $973$
1938$\ddagger$ $37.30$ $25.68$ $1364$ $46.51$ $25.46$ $177$ $35.93$ $25.44$ $1187$
Child mortality rate 1933$\ddagger$ $13.40$ $9.58$ $1140$ $15.54$ $7.11$ $167$ $13.03$ $9.90$ $973$
1938$\ddagger$ $16.81$ $8.73$ $1364$ $19.05$ $8.52$ $177$ $16.47$ $8.72$ $1187$
Panel C: Year Overall Treated Controlled
Labor supply Mean Std. Dev. Obs. Mean Std. Dev. Obs. Mean Std. Dev. Obs.
Labor force participation rate 1930$\ddagger$ $48.95$ $5.96$ $1140$ $45.76$ $6.10$ $167$ $49.50$ $5.76$ $973$
       Male 1930 $57.13$ $3.30$ $1140$ $56.90$ $3.61$ $167$ $57.17$ $3.24$ $973$
       Female 1930$\ddagger$ $40.77$ $10.65$ $1140$ $34.06$ $10.99$ $167$ $41.93$ $10.15$ $973$
Mining sector 1930$\ddagger$ $2.75$ $9.12$ $1140$ $17.13$ $17.64$ $167$ $0.28$ $1.63$ $973$
       Male 1930$\ddagger$ $3.16$ $10.09$ $1140$ $19.32$ $19.17$ $167$ $0.39$ $2.00$ $973$
       Female 1930$\ddagger$ $1.83$ $6.92$ $1140$ $11.85$ $14.30$ $167$ $0.11$ $0.99$ $973$
Agricultural sector 1930$\ddagger$ $65.82$ $23.27$ $1140$ $45.55$ $25.93$ $167$ $69.30$ $20.91$ $973$
       Male 1930$\ddagger$ $60.39$ $23.62$ $1140$ $38.84$ $25.01$ $167$ $64.09$ $21.29$ $973$
       Female 1930$\ddagger$ $74.03$ $21.90$ $1140$ $57.66$ $25.20$ $167$ $76.84$ $19.98$ $973$
Manufacturing sector 1930$\ddagger$ $10.85$ $7.95$ $1140$ $12.39$ $6.08$ $167$ $10.59$ $8.21$ $973$
       Male 1930$\ddagger$ $13.91$ $8.04$ $1140$ $16.80$ $7.12$ $167$ $13.41$ $8.08$ $973$
       Female 1930$\ddagger$ $5.93$ $8.78$ $1140$ $4.03$ $3.62$ $167$ $6.26$ $9.35$ $973$
Commerce sector 1930$\ddagger$ $8.51$ $7.13$ $1140$ $10.44$ $6.62$ $167$ $8.18$ $7.17$ $973$
       Male 1930$\dagger$ $7.86$ $6.44$ $1140$ $8.97$ $5.68$ $167$ $7.67$ $6.55$ $973$
       Female 1930$\ddagger$ $10.11$ $10.10$ $1140$ $14.28$ $10.64$ $167$ $9.40$ $9.83$ $973$
Domestic sector 1930$\ddagger$ $1.77$ $1.31$ $1140$ $2.05$ $1.05$ $167$ $1.72$ $1.34$ $973$
       Male 1930$\dagger$ $0.37$ $0.54$ $1140$ $0.27$ $0.25$ $167$ $0.38$ $0.57$ $973$
       Female 1930$\ddagger$ $4.21$ $3.93$ $1140$ $5.75$ $3.70$ $167$ $3.94$ $3.91$ $973$
Notes: $\ddagger$ and $\dagger$ in Panels A–C indicate that the mean difference between treatment and control groups is statistically significant at the 1% and 5% levels, respectively. The treatment (control) group is 0–5 (5-30) km from the centroid of a municipality that has a mine.
Panel A reports the summary statistics for the demographic outcomes: population, crude marriage rates, crude birth rates, and household size. The population is log-transformed in the regressions to deal with a skewed distribution in nature. Household size is the total number of people divided by the number of households. The crude marriage rate is the number of marriages per $1,000$ people. The crude birth rate is the number of live births per $1,000$ people. The marital fertility rate is the number of live births per $1,000$ married females. The crude death rate is the number of male (female) deaths per $1,000$ males (females).
Panel B reports the summary statistics for infant mortality, fetal death, and child mortality rates. The infant mortality rate is the number of infant deaths per $1,000$ live births. The fetal death rate is the number of fetal deaths per $1,000$ births. The child mortality rate is the number of deaths of children aged 1–4 per $1,000$ children aged 1–5. The variables with $\star$ have a set of censored observations; the results from the Tobit models are materially similar to the main results.
Panel C reports the summary statistics for labor force participation rates and employment shares. Labor force participation is the number of workers per $100$ people. The employment share is the number of workers in each sector per $100$ workers.
Sources: Data on the number of people and households are from the Statistics Bureau of the Cabinet (1935b, 1939). Data on the demographic variables are from the Statistics Bureau of the Cabinet (1932, 1938). Data on the labor force participation rates and employment shares are from the Statistics Bureau of the Cabinet (1933). Data on births, fetal, infant, and child deaths are from Aiikukai (1935); Social Welfare Bureau (1941).
C.3.1 Population, Marriage, Fertility, and Mortality
The official 1930 and 1935 Population Censuses published by the Statistics Bureau of the Cabinet document the number of people (by sex), households, and married females in municipalities. The number of marriages and live births used to calculate crude marriage and birth rates were obtained from the Municipal Vital Statistics of 1930 and 1935 published by the Statistics Bureau of the Cabinet. The data on the number of deaths by gender are from the same documents.
C.3.2 Early-life Mortality
Data on the number of fetal, infant, and child deaths were digitized using the official reports of municipal-level vital statistics of 1933 and 1938 published by the Aiikukai and Social Welfare Bureau. Data on the number of births and live births used to calculate fetal death and IMRs were obtained from the same report. Data on the number of children aged 1–5 were obtained from the 1930 Population Censuses.
C.3.3 Labor Supply
In Online Appendix D.4, the labor force participation rate, defined as the number of workers per $100$ people, is considered to investigate the impacts on local labor supplies. To analyze structural shifts, the employment share of workers in the mining, agricultural, manufacturing, commercial, and domestic sectors are used. These are calculated as the number of workers in each industry per $100$ workers. These variables are considered for both sexes to understand the potential gender bias in the shifts. These labor statistics were obtained from the prefectural part of the 1930 Population Census. [^54] The prefectural part comprises $47$ reports (one for each prefecture), which I collected and digitized. For simplicity, those reports are cited as one document, Statistics Bureau of the Cabinet (1935a). The full population census is conducted every ten years in Japan. The 1930 census surveyed all labor statistics. Thus, similar statistics were unavailable in the 1935 Population Census. The 1940 Population Census was also disorganized because of the Second World War. Panel C of Table C.4 presents the summary statistics of these variables by treatment status and census year. [^55] The employment share in the mining sector has a set of censored observations because mining is a localized economic activity. As a robustness check, it was confirmed that the estimates from the Tobit estimator Tobin (1958) are similar to the main results (not reported). Similar to the demographic outcomes, the mean differences are statistically significant for most variables.
C.4 Control Variables
C.4.1 Railway Stations
The location information of railway stations was obtained from the official shapefile file created by the Ministry of Land, Infrastructure, Transport and Tourism (https://nlftp.mlit.go.jp/ksj/gml/datalist/KsjTmplt-N05-v1˙3.html, accessed 29th July 2022). The railway stations used to match the 1931 (1936) mining dataset included all stations (i.e., national, public, and private rails) in the railway sections in 1931 (1936). Figure C.5a (C.5b) illustrates the distribution of stations in 1931 (1936). The shortest distance between the centroid of each municipality and the station was used as the distance from the nearest neighboring station. Both figures confirm that railway stations were not concentrated in the mining areas. This is consistent with evidence that accessibility to railways does not influence the results.
(a) 1931
(b) 1936
Figure C.5: Spatial Distribution of Railway Stations
Notes: The blue circles indicate the location of railway stations. The base map used in this figure is Figure 2. The treatment group highlighted in red includes the municipalities within 5 kilometers of a mine. The control group highlighted in pink includes the municipalities between 5 and 30 kilometers from a mine. The excluded municipalities are shown as empty lattices in the figures.
Source: Created by the author. The location data on the railway stations are from the Ministry of Land, Infrastructure, Transport and Tourism (website).
C.4.2 Seaports
Seaport locations were obtained from the official shapefile file created by the Ministry of Land, Infrastructure, Transport and Tourism (https://nlftp.mlit.go.jp/ksj/gmlold/datalist/gmlold˙KsjTmplt-C02.html, accessed 1st October 2022). The seaports used to match the 1931 (1936) mining dataset included all seaports (both commercial and fishery) listed in the 1931 (1936) edition of the Harbor Statistics of the Great Empire of Japan (Dainihon teikoku kōwan tōkei). The ports listed in the Harbor Statistics, but not included in the original shapefile of the Ministry of Land, Infrastructure, Transport, and Tourism, were added manually. Figure C.6a (C.6b) shows the distribution of seaports in 1931 (1936).
(a) 1931
(b) 1936
Figure C.6: Spatial Distribution of Sea Ports
Notes: The blue circles indicate the location of the seaports. The ports in Okinawa prefecture and a few ports in the islands far from the main island are not shown in the figures. The base map used in this figure is Figure 2. The treatment group highlighted in red includes the municipalities within 5 kilometers of a mine. The control group highlighted in pink includes the municipalities between 5 and 30 kilometers from a mine. The excluded municipalities are shown as empty lattices in the figures.
Source: Created by the author. The location data on the seaports are from the Ministry of Land, Infrastructure, Transport and Tourism (website).
C.4.3 Rivers
The location of rivers was obtained from the official shapefile provided by the Ministry of Land, Infrastructure, Transport, and Tourism (https://nlftp.mlit.go.jp/ksj/jpgis/datalist/KsjTmplt-W05.html, accessed August 21, 2018). The rivers used to match the mining dataset included those managed either by the government or prefectures. As the rivers had not moved, the location information is time constant. Figure C.7a (C.7b) shows the distribution of the rivers matched to the 1931 (1936) base map.
(a) 1931
(b) 1936
Figure C.7: Spatial Distribution of Rivers
Notes: The blue lines indicate the location of rivers. The rivers in Okinawa prefecture and a few points in the islands far from the main island are not shown in the figures. The base map used in this figure is Figure 2. The treatment group highlighted in red includes the municipalities within 5 kilometers of a mine. The control group highlighted in pink includes the municipalities between 5 and 30 kilometers from a mine. The excluded municipalities are shown as empty lattices in the figures.
Source: Created by the author. The location data of rivers are from the Ministry of Land, Infrastructure, Transport, and Tourism (website).
C.4.4 Elevation
The grid cell data on elevation are taken from the official GIS database of the Ministry of Land, Infrastructure, Transport, and Tourism (https://nlftp.mlit.go.jp/ksj/gml/datalist/KsjTmplt-G04-d.html, accessed February 24, 2023). The average elevation in a 1 km square grid cell is calculated using the elevation measured in the 250 square grid cells in a third mesh-code initially assigned by the Ministry of Land, Infrastructure, Transport, and Tourism. The nearest neighbor grid cell is then linked to each municipality. The average distance to the nearest neighbor is $0.49$km, which is sufficiently close to represent each municipality’s average elevation. Figure C.8 shows the spatial distribution of the elevation.
Figure C.8: Spatial Distribution of Elevation
Note: This figure illustrates the distribution of average elevation (in meters) in each municipality. The rivers in Okinawa prefecture and a few points in the islands far from the main island are not shown in the figures. The white circles indicate the location of coal mines in 1936. Source: Created by the author. The grid cell data on elevation are from the Ministry of Land, Infrastructure, Transport, and Tourism (website).
Table C.5: Summary Statistics: Control Variables
Year Coal mines GSC mines
Control Variables Mean Std. Dev. Obs. Mean Std. Dev. Obs.
Distance to station 1931 $5.38$ $5.84$ $1140$ $6.12$ $7.76$ $2498$
1936 $5.62$ $6.18$ $1364$ $5.56$ $6.92$ $3293$
Distance to port 1931 $26.22$ $22.19$ $1140$ $29.87$ $22.22$ $2498$
1936 $24.13$ $21.52$ $1364$ $32.66$ $24.28$ $3293$
Distance to river 1931 $0.99$ $1.59$ $1140$ $0.84$ $1.02$ $2498$
1936 $1.02$ $1.55$ $1364$ $0.90$ $1.93$ $3293$
Elevation 1931 $187.27$ $241.72$ $1140$ $182.97$ $207.41$ $2498$
1936 $181.30$ $233.42$ $1364$ $220.92$ $245.93$ $3293$
Notes: This table shows the summary statistics for the control variables by type of mines. GSC mines indicate gold, silver, and copper mines. Distances to the nearest station, port, and river are in kilometers. Elevation is in meters.
Sources: See Sections C.4.1 to C.4.4 for the data on the distance and elevation variables. Data on the location of mines are from the Association for Harmonious Cooperation (1932, 1937).
Appendix Appendix D Empirical Analysis Appendix
D.1 Attenuation due to Measurement Error
Consider a simplified projection of $y_{i}$ on $\textit{MineDeposit}_{i}$ as $y_{i}=\theta\textit{MineDeposit}_{i}+\varepsilon_{i}$. When $\textit{MineDeposit}_{i}$ has a classical measurement error, $\textit{MineDeposit}^{*}_{i}=\textit{MineDeposit}_{i}+u_{i}$, where $u_{i}$ satisfies mean zero property. As explained in the main text, $u_{i}$ can be regarded as the unobservable changes in the assignments between the measured years of outcome and treatment variables. Given the random nature of the location of mines (Section 5.1), the measurement error ($u_{i}$) is likely to be uncorrelated with the outcome, true treatment variable, and error in the true model. Under this setting, the observable model becomes $y_{i}=\theta^{*}\textit{MineDeposit}^{*}_{i}+\nu_{i}$, where $\nu_{i}=\varepsilon_{i}-\theta u_{i}$. Consequently, I use $\text{cov}(\textit{MineDeposit}^{*}_{i},\nu_{i})=E[\textit{MineDeposit}^{*}_{i}\nu_{i}]=-\theta\text{var}(u_{i})\neq 0$ for $\theta^{*}=\theta+\text{cov}(\textit{MineDeposit}^{*}_{i},\nu_{i})/\text{var}(\textit{MineDeposit}^{*}_{i})$.
D.2 Assumptions for the Identification
To simplify the discussion on the identification assumptions, I focus on the most straightforward structural equation, a dummy endogenous variable model taking the form:
$\footnotesize{\begin{split}y_{i}=\alpha+\beta\text{{MineDeposit}}_{i}+e_{i},\\ \text{{MineDeposit}}_{i}=\kappa+\tau\text{{Stratum}}_{i}+\epsilon_{i}.\end{split}}$ (20)
All the variables are defined in Section 5.1. Note that the same Greek symbols are used as those used in equations 2 and 3 to simplify the expressions of the model, albeit those shall be different parameters to those considered in Section 5.1.
Five assumptions are made to identify the parameter of interest under this setting: the stable unit treatment value assumption (hereafter SUTVA), random assignments, exclusion restriction, nonzero average causal effect of the IV, and monotonicity (Angrist et al. 1996). Section 5.1 shows that the assumptions of the random assignments and exclusion restriction are plausible given the nature of the spatial distributions of mines and stratum. This subsection then discusses the remaining three assumptions.
First, the SUTVA argues that the potential outcomes for municipality $i$ in the system (20) are only related to its own treatment status. Under my empirical setting that utilizes the municipal-level dataset, this assumption is less likely to be violated. This is because there are certain distances among different treated municipalities, even in the agglomerated area such as the Chikuhō coalfield. Moreover, to minimize the risk of SUTVA violation, a defined threshold is considered to define the treatment group. The results demonstrate that the baseline definition of the treatment distance is plausible to model the potential spillover effects in the regression (Online Appendix D.6).
Second, let $\textit{MineDeposit}_{i}(\textit{Stratum}_{i})$ be an indicator for whether $i$ is treated given the random assignments of the geological strata, under the SUTVA. The nonzero average causal effect indicates that $E[\textit{MineDeposit}_{i}(1)-\textit{MineDeposit}_{i}(0)]\neq 0$. This assumption clearly holds because the likelihood of having coal mines increases with the IV strata (Online Appendix C.2).
Third, the monotonicity assumption suggests that $\forall i$, $\textit{MineDeposit}_{i}(1)\geq\textit{MineDeposit}_{i}(0)$. In sum, a combination of the nonzero average causal effect and monotonicity assumptions argue that strong monotonicity should be satisfied at least in one unit as $\textit{MineDeposit}_{i}(1)>\textit{MineDeposit}_{i}(0)$ (Angrist et al. 1996). Although this assumption is an idea for the potential outcome and is untestable (Imbens and Angrist 1994), there is fundamentally no reason for mining companies to avoid the stratum that contains coal seams. Thus, the instrument should affect the selection decision monotonically.
D.3 Alternative Variance Estimator
To assess the potential influences of the local-scale spatial correlations, the standard errors clustered at the county level based on the cluster-robust covariance matrix estimator are used (Arellano 1987). The finite-sample adjustment proposed by Hansen (2007) is implemented in the estimator. Table D.1 summarizes the results based on the cluster-robust variance-covariance matrix estimator. All the results are materially similar to those of the baseline results in the main text.
Table D.1: Additional Results using the CRVE
Panel A: Population and Fertility Responses
Data: 1930 Census Data: 1935 Census
(1) Population (2) Crude (3) Marital (4) Population (5) Crude (6) Marital
    Marriage     Fertility     Marriage     Fertility
MineDeposit 0.6820.682***0.682 ∗ ∗ ∗ 2.732-2.732***- 2.732 ∗ ∗ ∗ 36.792-36.792***- 36.792 ∗ ∗ ∗ 0.7230.723**0.723 ∗ ∗ 1.482-1.482**- 1.482 ∗ ∗ $-9.734$
$(0.206)$ $(0.977)$ $(10.259)$ $(0.175)$ $(0.707)$ $(10.046)$
FEs and Controls Yes Yes Yes Yes Yes Yes
First-stage $F$-statistic $13.22$ $13.22$ $13.22$ $14.55$ $14.55$ $14.55$
Panel B: Mortality Changes
Data: 1930 Census Data: 1935 Census
(1) Male (2) Female (3) Male (4) Female
MineDeposit $0.698$ $-1.545$ $-0.578$ 2.010-2.010*- 2.010 ∗
$(1.470)$ $(1.591)$ $(1.168)$ $(1.217)$
FEs and Controls Yes Yes Yes Yes
First-stage $F$-statistic $13.22$ $13.22$ $14.55$ $14.55$
Panel C: Occupational Hazards and Pollutions
Analytical Sample
Full Treated Full Treated
(1) IMR (2) FDR (3) FDR (4) IMR (5) CMR (6) CMR
Panel C-1: 1933
MineDeposit 36.74036.740*36.740 ∗ 21.19421.194**21.194 ∗ ∗ $1.460$
$(22.064)$ $(8.929)$ $(4.436)$
Female Miner 0.0380.038***0.038 ∗ ∗ ∗ 0.1040.104***0.104 ∗ ∗ ∗ $0.004$
$(0.012)$ $(0.031)$ $(0.004)$
Male Miner 0.003-0.003*- 0.003 ∗ 0.010-0.010**- 0.010 ∗ ∗ $-0.001$
$(0.002)$ $(0.004)$ $(0.001)$
River accessibility $-2.122$ $0.367$ $-2.993$ $0.209$ 0.632-0.632**- 0.632 ∗ ∗ $0.603$
$(1.596)$ $(0.742)$ $(2.271)$ $(3.638)$ $(0.251)$ $(0.462)$
FEs and Controls Yes Yes Yes Yes Yes Yes
Estimator IV IV OLS OLS IV OLS
First-stage $F$-statistic $13.22$ $13.22$-- - --- - $13.22$-- -
Panel C-2: 1938
MineDeposit $18.013$ $6.436$ $1.967$
$(15.272)$ $(10.851)$ $(3.008)$
Female Miner 0.0330.033**0.033 ∗ ∗ 0.0730.073***0.073 ∗ ∗ ∗ 0.0070.007*0.007 ∗
$(0.013)$ $(0.028)$ $(0.004)$
Male Miner $-0.000$ $-0.001$ $-0.000$
$(0.001)$ $(0.002)$ $(0.001)$
River accessibility 2.164-2.164**- 2.164 ∗ ∗ $0.429$ $0.227$ $-0.145$ $-0.332$ $-0.148$
$(1.100)$ $(0.565)$ $(1.996)$ $(3.485)$ $(0.235)$ $(0.701)$
FEs and Controls Yes Yes Yes Yes Yes Yes
Estimator IV IV OLS OLS IV OLS
First-stage $F$-statistic $14.56$ $14.56$-- - --- - $14.56$-- -
Panel D: Social Changes
Data: 1930 Census Data: 1935 Census
(1) Sex Ratio (2) HH Size (3) Sex Ratio (4) HH Size
MineDeposit 0.1150.115***0.115 ∗ ∗ ∗ 0.534-0.534***- 0.534 ∗ ∗ ∗ 0.0650.065**0.065 ∗ ∗ 0.486-0.486**- 0.486 ∗ ∗
$(0.040)$ $(0.179)$ $(0.024)$ $(0.202)$
FEs and Controls Yes Yes Yes Yes
First-stage $F$-statistic $13.22$ $13.22$ $14.55$ $14.55$
***, **, and * represent statistical significance at the 1%, 5%, and 10% levels, respectively. Standard errors based on the cluster-robust covariance matrix estimator are reported in parentheses. Standard errors are clustered at the 107 (124) counties in the 1933 (1938) sample.
Notes: Panel A: The results for the population (log-transformed), crude marriage rate, and marital fertility rate are reported. Panel B: ‘Male’ and ‘Female’ indicate the male death rate and female death rate, respectively. Panels C-1 and C-2: The results for the 1933 and 1938 Vital Statistics samples are listed. Columns 1–2 show the results for the entire sample. Columns 3–4 show the results for the municipalities with coal mines (i.e., the municipalities included in the treatment group). Columns 5–6 show the results for the full sample and the municipalities with coal mines, respectively. The IMR, FDR, and CDR indicate the infant mortality, fetal death, and child mortality rates, respectively. Panel D: ‘Sex Ratio’ and ‘HH Size’ indicate the sex ratio (male/female) and average household size, respectively. Table C.4 lists summary statistics on the dependent variables. All the estimates listed in Panels A, B, and D are from the IV estimator. All the regressions include the city and town fixed effects, railway accessibility, port accessibility, river accessibility, and elevation.
D.4 Gender-Biased Structure Shift: Evidence from the 1930 Census
Table D.2: Labor Force Participation and Structural Shifts
Panel A: Male Workers
Dependent Variable
Employment Share
(1) LFPR (2) Mining (3) Agricultural (4) Manufacturing (5) Commerce (6) Domestic
MineDeposit $-0.300$ 33.94033.940***33.940 ∗ ∗ ∗ 31.984-31.984***- 31.984 ∗ ∗ ∗ $1.526$ 2.777-2.777*- 2.777 ∗ 0.381-0.381***- 0.381 ∗ ∗ ∗
$(1.157)$ $(4.852)$ $(5.894)$ $(2.014)$ $(1.470)$ $(0.132)$
City and Town FEs Yes Yes Yes Yes Yes Yes
Railway accessibility Yes Yes Yes Yes Yes Yes
Port accessibility Yes Yes Yes Yes Yes Yes
River accessibility Yes Yes Yes Yes Yes Yes
Elevation Yes Yes Yes Yes Yes Yes
Observations 1,140 1,140 1,140 1,140 1,140 1,140
Estimator IV IV IV IV IV IV
First-stage $F$-statistic $36.05$ $36.05$ $36.05$ $36.05$ $36.05$ $36.05$
Mean of the DV $57.13$ $3.16$ $60.39$ $13.91$ $7.86$ $0.37$
Std. Dev. of the DV $3.30$ $10.09$ $23.62$ $8.04$ $6.44$ $0.54$
Panel B: Female Workers
Dependent Variable
Employment Share
(1) LFPR (2) Mining (3) Agricultural (4) Manufacturing (5) Commerce (6) Domestic
MineDeposit 16.308-16.308***- 16.308 ∗ ∗ ∗ 24.36424.364***24.364 ∗ ∗ ∗ 22.005-22.005***- 22.005 ∗ ∗ ∗ 10.341-10.341***- 10.341 ∗ ∗ ∗ $1.777$ 4.2014.201***4.201 ∗ ∗ ∗
$(3.644)$ $(4.049)$ $(5.525)$ $(2.134)$ $(2.516)$ $(1.312)$
City and Town FEs Yes Yes Yes Yes Yes Yes
Railway accessibility Yes Yes Yes Yes Yes Yes
Port accessibility Yes Yes Yes Yes Yes Yes
River accessibility Yes Yes Yes Yes Yes Yes
Elevation Yes Yes Yes Yes Yes Yes
Observations 1,140 1,140 1,140 1,140 1,140 1,140
Estimator IV IV IV IV IV IV
First-stage $F$-statistic $36.05$ $36.05$ $36.05$ $36.05$ $36.05$ $36.05$
Mean of the DV $40.77$ $1.83$ $74.03$ $5.93$ $10.11$ $4.21$
Std. Dev. of the DV $10.65$ $6.92$ $21.90$ $8.78$ $10.10$ $3.93$
***, **, and * represent statistical significance at the 1%, 5%, and 10% levels, respectively. Standard errors based on the heteroskedasticity-robust covariance matrix estimator are reported in parentheses.
Notes: Panels A and B show the results for male and female worker samples, respectively, from the 1930 Population Census (Panel C of Table C.4). Column 1 shows the results for the labor force participation rates (%), whereas Columns 2–6 show the results for employment share in each industrial sector (%).
In this section, I aim to evaluate how the coal mines influence the local industrial structure using labor supply statistics from the 1930 population census. Theory predicts the positive correlation between the resource extractions and local labor supply. Allcott and Keniston (2018) provides evidence of such a relationship for the oil and gas extraction industries in the post-World War II United States (U.S.) economy. My results show that the impacts of the coal mine on local labor supply depended on gender. While the local labor supply for male workers was not influenced by the coal extractions, that for female workers decreased with a meaningful magnitude. The rapid mechanization and institutional change observed in interwar Japan, which is not considered in the case of the postwar U.S. economy, may explain the gap in the findings. The labor-saving technological advances might had reduced the labor required in the mining sector and the labor regulations had further displaced women from the sector.
Male Workers
Panel A of Table D.2 presents the results for the male workers. Column 1 shows that the estimate for the male LFP is very close to zero and statistically insignificant. This indicates that coal mines did not influence the labor supply of male workers in the local economy. Columns 2–6 show the results for the employment share of male workers. Column 2 indicates that coal mines increased the mining sector’s employment share by approximately 34%, leading to a similar decline in the agricultural sector’s employment share (column 3). This suggests that the structural shift in male workers’ employment occurred mainly in the agricultural sector. In column 4, the estimate for the manufacturing sector is positive, suggesting a marginal spillover effect of mines on the manufacturing industry in the mining areas. However, this effect is not statistically significant. Compared with the case of the 1970s U.S., [^56] Black et al. (2005) found that the 1970s coal boom in the U.S. increased employment and earnings with modest spillovers into the non-mining sectors. the rapid mechanization may not have created the time window for generating spillover effects in the case of interwar Japan. The estimates for the commercial and domestic sectors are moderately negative and close to zero (columns 5–6, respectively). This is consistent with male workers being less likely to work in these service sectors (Panel C of Table C.4).
Female Workers
The results for female labor listed in Panel B demonstrate different responses. Column 1 shows a statistically significantly negative estimate, providing evidence that the coal mines decreased the female LFP. The estimate suggests that coal mines decreased the female LFP rate by approximately $16$%. This may be partly explained by the fact that the gender wage gaps in mining sector were relatively lower than those in the surrounding agrarian area (Panel B of Table 1). [^57] The clear gender difference in the labor force participation appears to be consistent with the wealth effect (“spending effect” in the Dutch-Disease literature) of mines. It may result in higher wage rates and lower overall non-resource GDP (Caselli and Coleman II 2001). Although it is difficult to analyze the impact of mines on wages, the results align with the findings on structural shifts presented in this subsection. Columns 2–6 show the results for the employment share of female workers. Column 2 indicates that coal mines increase the mining sector’s employment share by $24$% and decrease the agricultural sector’s share by the same degree (column 3). The estimated magnitude is smaller than that for male workers (Panel A). This is logical because, on average, females were less likely to work in mines than males (Panel A of Table 1). While the manufacturing sector’s employment share decreases in the mining area (column 3), the estimates for the commercial and domestic sectors are moderately positive (columns 5–6, respectively). This appears to be in line with the increased relative demand for personal services from mining workers in mining areas (Caselli and Coleman II 2001).
D.5 Conditional Expectation of the Truncated Normal Distribution
Let $x\sim\mathcal{N}(\mu,\,\sigma^{2})$ be the initial health endowment of a fetus in utero and $x_{s}$ be a survival threshold. The conditional expectation of the truncated normal distribution can be derived as:
[…]
This implies that while the “scarring” mechanism ($\mu^{*}<\mu$) shifts the conditional mean toward the left, the “selection” mechanism ($x_{s}^{*}>x_{s}$) shifts the conditional mean toward the right.
D.6 Testing the Validity of Threshold
(a) Population
(b) Fertility
(c) Mortality
(d) IMR
(e) Sex Ratio
(f) Household Size
Figure D.1: Heterogeneity in the Impact of Mines by Distance from the Mines
Notes: This figure shows the results from the expanded regressions of equation 2 that includes the five indicator variables for each 5km bin from a mine to 25 km away from the origin. 0–5 km distance from mines indicates the municipalities within 5 km of a mine. 5–10 (10–15, 15–20, 20–25) km distance from mines indicates the municipalities within $(5,10]$ ($(10,15],(15,20],(20,25]$) km of a mine. (25–30] km bin is used as a reference group.
Figures D.1aD.1bD.1cD.1dD.1e, and D.1f show the results for the local population, marital fertility rate, female mortality rate, infant mortality rate, sex ratio, and average household size, respectively. The dots and solid lines with caps show the estimates and their 95 percent confidence intervals, respectively. The confidence intervals are calculated using the standard errors based on the heteroskedasticity-robust covariance matrix estimator.
Source: Created by the author.
The validity of the threshold for the exposure variable (MineDeposit in equation 2) is assessed using the main outcome variables. Figure D.1 provides the results from the expanded regression of equation 2 that includes the five indicator variables for each 5 km bin from a mine. Figures D.1aD.1bD.1cD.1dD.1e, and D.1f show the results for the local population, marital fertility rate, female mortality rate, infant mortality rate, sex ratio, and average household size, respectively. Note that these estimates obtained from the OLS estimator are systematically attenuated toward zero. This means that the estimates for the dependent variables with relatively moderate effects based on the IV estimator, such as fertility and mortality, are flatter. In addition, the estimates shown in these figures are the relative values to the reference group, meaning that the statistical significance does not provide a meaningful interpretation of these figures. Therefore, the important point herein is that the shape of the estimates for the dependent variables that captured clear effects in Section 5 does show a clear decreasing trend concerning the distance from mine. Therefore, the results for the population (Figure D.1a) and sex ratio (Figure D.1e) are useful for the purpose of this test. [^58] In other words, the other variables, such as fertility and mortality, are not useful herein because the main results do not have systematic effects. For example, the estimate for 1935 marital fertility is statistically insignificant in the main text (Table 3). This means that the estimates outside the 5km threshold can be similar to the estimate for the 0-5km bin. Figure D.1b confirms this relationship. Overall, the estimates are larger in the absolute sense within $5$ km, whereas those for the areas outside the 5 km bin are close to zero in Figures D.1a and D.1e. This evidence provides a basis for defining the exposure variable that uses $5$ km as the threshold.
A 10-15 km bin for the IMR in 1938 shown in Figure D.1d shows a clear hike compared with the other estimates for the bins outside the 5km threshold. Since the main estimate for the 1938 IMR is not statistically significant (Table 5), this may be useful to analyze the reasons behind this hike. To detect the region causing the hike, the entire sample is stratified into two subsamples: the municipalities belonging to Fukuoka prefecture and the others. The former might be an influential prefecture in the entire sample because it includes a number of mines in the Chikuhō coalfield. Figure D.2 shows the estimated coefficients for each subsample, confirming that the hike originated from the Fukuoka prefecture variations. Next, Figure D.3a illustrates the distribution of the bins in Fukuoka, and Figure D.3b clips the 10-15 km bin in the map. From a comparison of the distribution of the bins and the IMRs shown in Figure D.3c, it appears that some municipalities in the 10-15 km bin suffered a relatively higher risk of infant deaths, which is consistent with the tendency in Figure D.2. Note that, given a point of origin (i.e., a mine), a point on the concentric circle is a function of the distance. Therefore, the hike in the 10-15 km bin is considered to be a random event, given that the location of mines is at least conditionally independent as described in Section 5. Another evidence supporting this randomness is that such a hike in the 10-15km bin is not observed in the results for the 1933 sample (see Figure D.1d). If there were systematic factors, both the 1933 and 1938 samples likely suffered the more significant risks of infant deaths in this bin. Similarly, the fact that only a 10-15 bin in Fukuoka prefecture suffers such a hike supports this argument. Overall, the hike is negligible because any systematic endogenous mechanism in the regression does not generate it.
To summarize, the hike observed in the 10-15 km bin shall be randomly occurred, but may not be led by any systematic factors that relate to the distribution of mines. [^59] Note that the models considered herein still allow the existence of unobservables that relate only to the IMR. These include the social norms for the child-rearing because these socioeconomic factors unrelated to the distribution of mines. Although a large typhoon hit Miyazaki and Kagoshima prefectures in 1938, which might have increased the mortality rates, both prefectures are not included in the dataset. Given the empirical setting, this also makes sense because using many bins reduces the available observations in each bin, making the moment estimator vulnerable to this type of random shock. [^60] As shown in Figure D.5, the infant mortality rates in Japan had moderate fluctuations during the 1930s, and the IMR in 1938 had slightly higher values than the surrounding years. This may further make the cross-sectional variation in the mortality rates greater. As for the main analysis in Column (1) in Panel B of Table 5, this sort of hike in a specific bin should not be an influential event because the observations in the 10-15km bins, particularly of Fukuoka, become a small proportion of the control group.
Figure D.2: Heterogeneity in the Impact of Mines by Distance from the Mines: The IMRs in 1938 in Fukuoka and Other Prefectures
Notes: This figure shows the results from the expanded regressions of equation 2 that include the five indicator variables for each 5km bin from a mine to 25 km away from the origin. 0–5 km distance from mines indicates the municipalities within 5 km of a mine. 5–10 (10–15, 15–20, 20–25) km distance from mines indicates the municipalities within $(5,10]$ ($(10,15],(15,20],(20,25]$) km of a mine. (25–30] km bin is used as a reference group.
Figure D.2 summarizes the results for the infant mortality rate in Fukuoka prefecture and the other prefectures. The dots and solid lines with caps show the estimates and their 95 percent confidence intervals, respectively. The confidence intervals are calculated using the standard errors based on the heteroskedasticity-robust covariance matrix estimator.
Source: Created by the author.
(a) Classification
(b) 10-15 km Bins
(c) IMR in 1938
Figure D.3: Spatial Distribution of Treatment Bins and Infant Mortality Rate
in Fukuoka Prefecture
Notes: Figure D.3a shows the spatial distribution of all the treatment bins used in Figure D.2. Figure D.3b highlights the location of the 10-15 km bin in Figure D.3a. Figure D.3c shows the infant mortality rates in 1938.
Source: Created by the author.
D.7 Sensitivity to the Regional Heterogeneity in Endowment of Male Miners
I assess the influence of regional heterogeneity in female labor patterns. Nishinarita (1985) argues that, due to its geographical characteristics, coal mines in Hokkaidō (a northernmost island) tended to employ men from rural villages in the Tōhoku (northeastern) region of the main island as their main miners (see also Ogino 1993, p. 32). This forced the coal mines in Hokkaidō to use a sloping-face haulage method instead of using female miners (ato-yama). As a result, female miners in the Hokkaidō mines were mainly engaged in the coal selection process rather than working in pits. It is unlikely that this regional heterogeneity could affect the location of mines because the location of coal mines depends on the presence or absence of coal-bearing strata (Section 5.1). In addition, the number of mines in Hokkaidō is relatively small, meaning that such regional heterogeneity in labor supply is less likely to disturb the main estimates. Despite this, to test whether such features lead to omitted variable bias, an indicator variable for the municipalities of the Hokkaidō prefecture was added in the regressions. Table D.3 shows the results from the expanded regressions. The results are materially similar to those of the main results presented in the main text.
Table D.3: Additional Results: IV Regressions including a Hokkaidō Dummy
Panel A: Population and Fertility Responses
Data: 1930 Census Data: 1935 Census
(1) Population (2) Crude (3) Marital (4) Population (5) Crude (6) Marital
    Marriage     Fertility     Marriage     Fertility
MineDeposit 0.7990.799***0.799 ∗ ∗ ∗ 2.897-2.897***- 2.897 ∗ ∗ ∗ 26.370-26.370**- 26.370 ∗ ∗ 0.8140.814***0.814 ∗ ∗ ∗ 1.620-1.620**- 1.620 ∗ ∗ $-3.173$
$(0.195)$ $(0.786)$ $(9.271)$ $(0.177)$ $(0.635)$ $(7.993)$
FEs and Controls Yes Yes Yes Yes Yes Yes
First-stage $F$-statistic $37.86$ $37.86$ $37.86$ $47.57$ $47.57$ $47.57$
Panel B: Mortality Changes
Data: 1930 Census Data: 1935 Census
(1) Male (2) Female (3) Male (4) Female
MineDeposit $0.295$ $-1.885$ $-0.781$ 2.291-2.291**- 2.291 ∗ ∗
$(1.400)$ $(1.362)$ $(1.139)$ $(1.057)$
FEs and Controls Yes Yes Yes Yes
First-stage $F$-statistic $37.86$ $37.86$ $47.57$ $47.57$
Panel C: Occupational Hazards and Pollutions
Analytical Sample
Full Treated Full Treated
(1) IMR (2) FDR (3) FDR (4) IMR (5) CMR (6) CMR
Panel C-1: 1933
MineDeposit 40.39240.392***40.392 ∗ ∗ ∗ 19.92519.925**19.925 ∗ ∗ $3.058$
$(15.620)$ $(9.396)$ $(2.767)$
Female Miner 0.0300.030**0.030 ∗ ∗ 0.0970.097***0.097 ∗ ∗ ∗ 0.0070.007*0.007 ∗
$(0.012)$ $(0.031)$ $(0.004)$
Male Miner $-0.002$ 0.009-0.009**- 0.009 ∗ ∗ 0.001-0.001*- 0.001 ∗
$(0.002)$ $(0.004)$ $(0.000)$
River accessibility 2.655-2.655**- 2.655 ∗ ∗ $0.552$ $-2.008$ $0.894$ 0.866-0.866***- 0.866 ∗ ∗ ∗ $0.365$
$(1.052)$ $(0.698)$ $(2.241)$ $(3.830)$ $(0.226)$ $(0.469)$
FEs and Controls Yes Yes Yes Yes Yes Yes
Estimator IV IV OLS OLS IV OLS
First-stage $F$-statistic $37.86$ $37.86$-- - --- - $37.86$-- -
Panel C-2: 1938
MineDeposit $18.253$ $6.090$ $2.229$
$(11.714)$ $(7.762)$ $(2.397)$
Female Miner 0.0310.031**0.031 ∗ ∗ 0.0710.071**0.071 ∗ ∗ 0.0070.007*0.007 ∗
$(0.014)$ $(0.028)$ $(0.004)$
Male Miner $-0.000$ $-0.001$ $-0.000$
$(0.001)$ $(0.002)$ $(0.001)$
River accessibility 2.199-2.199**- 2.199 ∗ ∗ $0.480$ $0.506$ $0.197$ 0.370-0.370*- 0.370 ∗ $-0.140$
$(1.033)$ $(0.588)$ $(2.019)$ $(3.452)$ $(0.216)$ $(0.718)$
FEs and Controls Yes Yes Yes Yes Yes Yes
Estimator IV IV OLS OLS IV OLS
First-stage $F$-statistic $47.57$ $47.57$-- - --- - $47.57$-- -
Panel D: Social Changes
Data: 1930 Census Data: 1935 Census
(1) Sex Ratio (2) HH Size (3) Sex Ratio (4) HH Size
MineDeposit 0.1250.125***0.125 ∗ ∗ ∗ 0.443-0.443***- 0.443 ∗ ∗ ∗ 0.0740.074**0.074 ∗ ∗ 0.408-0.408***- 0.408 ∗ ∗ ∗
$(0.034)$ $(0.131)$ $(0.022)$ $(0.129)$
FEs and Controls Yes Yes Yes Yes
First-stage $F$-statistic $37.86$ $37.86$ $47.57$ $47.57$
***, **, and * represent statistical significance at the 1%, 5%, and 10% levels, respectively. Standard errors based on the heteroskedasticity-robust covariance matrix estimator are reported in parentheses.
Notes: Panel A: The results for the population (log-transformed), crude marriage rate, and marital fertility rate are reported. Panel B: ‘Male’ and ‘Female’ indicate the male death rate and female death rate, respectively. Panels C-1 and C-2: The results for the 1933 and 1938 Vital Statistics samples are listed. Columns 1–2 show the results for the entire sample. Columns 3–4 show the results for the municipalities with coal mines (i.e., the municipalities included in the treatment group). Columns 5–6 show the results for the full sample and the municipalities with coal mines, respectively. The IMR, FDR, and CDR indicate the infant, fetal, and child mortality rates, respectively. Panel D: ‘Sex Ratio’ and ‘HH Size’ indicate the sex ratio (male/female) and average household size, respectively. Table C.4 lists summary statistics on the dependent variables. All the estimates listed in Panels A, B, and D are from the IV estimator. All the regressions include the city and town fixed effects, railway accessibility, port accessibility, river accessibility, elevation, and a Hokkaidō dummy.
D.8 Alternative Identification Strategy
The IV estimation was used to obtain the baseline estimates. For the outcome variables that can be used as a panel dataset, an alternative estimation strategy could be the DID approach, which uses both cross-sectional and within variations for the identification. To implement the DID technique, municipalities that had been treated throughout the measured years would have to be excluded. Simultaneously, municipalities that had newer mines during the sampled periods were left in. However, the number of municipalities with coal mines increased by only 23 between 1931 and 1936 (Online Appendix A.4 presents the raw number of mines). Consequently, as explained in Section 4.1 in detail, the proportion of treated municipalities decreased from $15$% to $13$% (Panel A of Table C.3). This feature of the assignments makes it difficult to employ the DID approach as the main identification strategy. This is because the information that can be used for the identification is insufficient. [^61] Similarly, the number of municipalities with GSC mines increased by only 55 mines from 57 to 112 between 1931 and 1936 (see Online Appendix A.4 for the number of mines). The proportion of treated municipalities remained almost unchanged, from $5.0$% to $5.3$% (Panel A of Table C.3).
Nevertheless, I provide a brief case study using the DID approach for the subsample, including Ogawa coal mine in Iwate prefecture and the Miyoshiyama coal mine in Akita prefecture. As indicated in Figure 2b, both mines were located in the northeastern (Tōhoku) region, where the coal mines did not exist in 1931. From this perspective, it is expected to provide relatively clean experimental conditions for the DID estimation, although the available sample size is limited.
For municipal $i$ at time $t\in\{1930,1935\}$, the two-way fixed effect (TWFE) model for the DID estimation is specified as follows:
$\footnotesize{\begin{split}\textit{y}_{i,t}=\pi+\delta\textit{D}_{i,t}+\mathbf{x}^{\prime}_{i,t}\bm{\phi}+\eta_{i}+\lambda_{t}+\omega_{i,t},\end{split}}$ (21)
where $\eta$, $\lambda$, and $\omega$ are the municipal fixed effect, year fixed effect, and random error term, respectively. $\textit{D}_{i,t}$ is a treatment indicator variable that takes one for the treatment group in the post-treatment period and zero for otherwise. $\mathbf{x}$ includes the accessibility measures on the stations and sea ports. Accessibility measure on the rivers is not included as it is a time-constant variable. However, the municipal fixed effect effectively controls for all these time-constant unobservable factors, including the geological strata and terrain that were potentially correlated with the location of mines.
It should be noted that this is a two-by-two DID setting for the TWFE model. It provides the unbiased DID estimator ($\hat{\delta}$) for the average treatment effect in the treated municipalities in the post-treatment period (Abadie 2005). [^62] For a comprehensive review on the recent discussions on the TWFE-DID models, see de Chaisemartin and D’Haultfœuille (2022). The identification assumptions are the independence of the treatment, which has already been confirmed, and the parallel trends assumption. Figure D.4 shows the time-series plots of the outcome variables used in the DID analysis for the Iwate and Akita subsample in Table D.4. Although there are moderate fluctuations due to the small sample size in each group, these outcomes do not show peculiar trends but show materially similar trends in the pretreatment years (i.e., 1925 and 1930). The outcomes of the treatment group in the post-treatment period (i.e., 1935) show clear kinks in most cases, whereas those for the control group show marginal changes in the post-treatment period. A similar check for the mortality rates is unavailable because the data on deaths in 1925 are unavailable.
(a) Population
(b) Marriage
(c) Fertility
(d) Sex Ratio
(e) Household Size
Figure D.4: Trends in the Two Mining Areas in Iwate and Akita
Notes: This figure shows the time-series plots of the outcomes of the treatment and control groups in Iwate and Akita prefectures used in the regressions shown in Table D.4. Figures D.4aD.4bD.4cD.4e, and D.4d show the average number of people, average crude marriage rate, average marital fertility rate, average sex ratio and average household size, respectively. The data on the mortality in 1925 are unavailable.
Source: Created by the author.
The cluster-robust variance-covariance matrix estimator is used to deal with the heterosckedasticity and serial dependency (Arellano 1987).
Table D.4: Additional Results: Difference-in-Differences
for Two Coal Mine Areas in Iwate and Akita Prefectures
Panel A: Population, Fertility Responses, and Mortality Changes
Dependent Variable
(1) Population (2) Marriage (3) Fertility (4) Mortality (Male) (5) Mortality (Female)
MineDeposit 0.0760.076***0.076 ∗ ∗ ∗ $-1.482$ 22.39322.393***22.393 ∗ ∗ ∗ $1.675$ $0.013$
$(0.023)$ $(1.510)$ $(7.809)$ $(2.386)$ $(2.377)$
Municipal and Year FEs Yes Yes Yes Yes Yes
Railway accessibility Yes Yes Yes Yes Yes
Port accessibility Yes Yes Yes Yes Yes
Observations 136 136 136 136 136
Clusters 68 68 68 68 68
Years 1930 and 35 1930 and 35 1930 and 35 1930 and 35 1930 and 35
Estimator DID DID DID DID DID
Panel B: Occupational Hazards and Social Changes
Dependent Variable
(1) IMR (2) FDR (3) Child MR (4) Sex Ratio (5) HH Size
MineDeposit $22.716$ $9.174$ $1.144$ 0.0450.045***0.045 ∗ ∗ ∗ 0.103-0.103**- 0.103 ∗ ∗
$(24.769)$ $(10.866)$ $(5.197)$ $(0.011)$ $(0.044)$
Municipal and Year FEs Yes Yes Yes Yes Yes
Railway accessibility Yes Yes Yes Yes Yes
Port accessibility Yes Yes Yes Yes Yes
Observations 136 136 136 136 Yes
Clusters 68 68 68 68 Yes
Years 1930 and 35 1930 and 35 1930 and 35 1930 and 35 1930 and 35
Estimator DID DID DID DID DID
***, **, and * represent statistical significance at the 1%, 5%, and 10% levels, respectively. Standard errors based on the cluster-robust covariance matrix estimator are reported in parentheses.
Notes: This table shows the results for the coal mines subsample in Iwate and Akita prefectures. Panel A: Column 1 shows the result for the log-transformed population. Columns 2–5 show the results for the crude marriage rate, marital fertility rate, male mortality rate, and female mortality rate, respectively. Panel B: Columns 1-3 show the results for the infant mortality, fetal death, and child mortality rates, respectively. Columns 4–5 show the results for the sex ratio and average household size, respectively.
Table D.4 provides the results. In Panel A, column 1 shows the result for the local population, whereas columns 2–5 show those for the marriage rate, marital fertility rate, and mortality rates. In Panel B, columns 1-3 present the results for the early-life mortality rates, whereas columns 4–5 show the results for the sex ratio and household size, respectively. Overall, the DID results align with the findings from the baseline results. First, the opening of these two coal mines increased the local population by approximately $8$%. Although the estimated magnitude is much smaller than the baseline magnitude listed in Table 2636363Notice that the estimates are compared with those for the results in 1935 or 1938 because the DID estimator captures the average treatment effect in the treated municipalities at the post-treatment period., this is consistent with the fact that both Ogawa and Miyoshiyama coal mines are small-scale mines with approximately $100$ miners. [^64] In fact, this magnitude is rather similar to those estimated for the smaller coal mines in my expanded regressions (Table D.6). The estimates for fertility and mortality rates are also smaller than the baseline estimates, albeit the fertility pathway is still observed in the statistical sense. This aligns with the small influence of the small-scale mines (Online Appendix D.11). Moreover, both coal mines used few female miners: only $9$ and $44$ miners in Ogawa and Miyoshiyama coal mines, respectively. Thus, the regulations were less likely to influence these mines. To summarize, the local population growth in both mining areas was mainly led by the influx of male workers, which is supported by the estimate for the sex ratio and average household size (columns 4–5).
D.9 Alternative Minerals: Heavy Metal Mines
Table D.5: Additional Results: Heavy Metal Mines
Panel A: 1930 sample
Dependent Variable
(1) Population (2) Fertility (3) MR (Male) (4) MR (Female) (5) SR (6) Size
MineDeposit 0.1740.174***0.174 ∗ ∗ ∗ $-0.080$ $0.192$ 0.763-0.763**- 0.763 ∗ ∗ $0.004$ 0.152-0.152***- 0.152 ∗ ∗ ∗
$(0.049)$ $(3.051)$ $(0.401)$ $(0.376)$ $(0.007)$ $(0.053)$
FEs and Controls Yes Yes Yes Yes Yes Yes
Observations 2498 2498 2498 2498 2498 2498
Estimator OLS OLS OLS OLS OLS OLS
Panel B: 1935 sample
Dependent Variable
(1) Population (2) Fertility (3) MR (Male) (4) MR (Female) (5) SR (6) Size
MineDeposit 0.2480.248***0.248 ∗ ∗ ∗ $0.628$ $-0.593$ $-0.553$ 0.0330.033***0.033 ∗ ∗ ∗ $-0.005$
$(0.049)$ $(2.185)$ $(0.370)$ $(0.387)$ $(0.008)$ $(0.049)$
FEs and Controls Yes Yes Yes Yes Yes Yes
Observations 3293 3293 3293 3293 3293 3293
Estimator OLS OLS OLS OLS OLS OLS
***, **, and * represent statistical significance at the 1%, 5%, and 10% levels, respectively. Standard errors based on the heteroskedasticity-robust covariance matrix estimator are reported in parentheses.
Notes: This table shows the results for the GSC mines sample. Panels A and B show the results for the 1930 and 1935 samples, respectively. Column 1 shows the results for the log-transformed population. Column 2 shows the results for the marital fertility rate. Columns 3 and 4 show the results for the male and female mortality rates, respectively. Columns 5 and 6 show the results for the sex ratio and average household size, respectively.
The validity of the main results on coal mines is assessed by comparing the results for heavy metal mines. Specifically, I focus on three heavy metals: gold, silver, and copper (GSC). This is because GSC mines comprise more than $80$% of all metal mines and generally extract these metals in one place. Although the AHC documented other types of metal mines, they were much smaller subsets: the number of lead and zinc, pig iron and steel, and tin mines recorded were only one, five, and four in 1931, respectively.
Historical evidence indicates that coal mining requires many more workers than heavy metal mining because coal is a bulky mineral (Section 2). Table C.3 indicates that the scale of GSC mines was much smaller than that of coal mines. In addition, unlike coal mines, the GSC mines were scattered across the archipelago and were less likely to agglomerate than coal mines (Figure C.2). The “enclave effect” suggests that some GSC mines are less likely to interact with local economies (Aragón and Rud 2013, 2015). [^65] A potentially relevant study is Stijns (2005); using a country-level dataset, the author provided evidence suggesting that the impact of resource abundance on socioeconomic outcomes may vary among resource types. This means that the impact of GSC mines on the regional population growth must be smaller than that of coal mines. To assess this, the estimates of the reduced-form equation 2 for the coal and GSC mines are compared because, unlike coal and fuel minerals, no specific stratum can predict the spatial distribution of GSC mines.
Table D.5 lists the results. Panel A (B) summarizes the results for the 1930 (1935) sample. The estimates for the local population listed in column 1 are $0.174$ for 1930 sample and $0.248$ for 1935 sample, respectively. These are less than half of the estimates for the coal mine sample (column 3 of Panels A and B in Table 2). This result confirms the baseline findings on the coal mines.
However, the results for the other demographic variables differ greatly from those for the coal mines. The estimates for the fertility and mortality rates are close to zero and statistically insignificant in most cases (columns 2–4). The estimate for female morality in 1930 is statistically significantly negative. While this might capture the wealth effects of the GSC mines, the estimate becomes less clear in 1935. Similarly, columns 5 and 6 do not provide systematic results on the sex ratio and average household size in the GSC mining area. These results imply that the regional growth in the GSC mining area may be led by mechanisms different from those for the coal mines. This reflects the fundamental difference between coal and heavy metal mines as materials.
D.10 Time-Series Plots of the IMR and FDR
Figure D.5 shows the national average IMR (solid line) and FDRs (dashed line). Both series show secular decreasing trends between 1925 and 1940.
Figure D.5: The Infant Mortality and Fetal Death Rates
between 1925 and 1940
Note: This figure shows the national average infant mortality (IMR; solid line) and fetal death rates (FDR; dashed line) in Japan between 1925 and 1940. The IMR is the number of infant deaths per 1,000 live births. The FDR is defined as the number of fetal deaths per 1,000 births.
Sources: Created by the author. Data on the number of infant and fetal deaths are from the Statistics Bureau of the Cabinet (1926–1942).
D.11 Heterogeneity with Respect to Scale of Mines
While most of the municipalities with mines have a similar number of mines (i.e., one or two mines), several municipalities have many mines. [^66] In the 1930 (1935) sample, 85 (108) out of 93 (115) municipalities with mines have one to four mines, whereas the other 8 (7) municipalities have more than five mines. The average number of miners in the former group was approximately 1,300 (1,400) in 1930 (1935) sample, whereas that in the latter group was approximately 4,000 (5,400) in 1930 (1935) sample. Essentially, causal effects are heterogeneous across units. In my setting, this means that the municipalities with large-scale mine could have more significant impact than those with small-scale mines. Estimating the heterogeneous treatment effects in detail is beyond the scope of this study. Estimating such varying effects also presents a practical issue. Stratifying the treatment group disturbs the inference due to the small sample size in each subsample. Moreover, deciding a certain threshold for the scale of mine is essentially impossible. Despite this, it may be useful to determine whether the potential heterogeneity in the treatment effect concerning the scale of mines is reasonable. To assess the heterogeneity, the treatment group is first stratified into two subgroups using the median of the number of miners as a threshold. The key treatment indicator variable in equation 2 is then replaced with the two indicator variables for the two subgroups.
Table D.6 presents the results. Panels A and B show the results for the 1930 and 1935 samples, respectively. Columns 1–6 list the estimates for the local population, marital fertility, mortality rates, sex ratio, and average household size. Column 1 suggests that the relatively large-scale mines had more significant impacts on the local population’s growth. Fertility responses and mortality changes are clearly observed in the large-scale mines (columns 2 and 3). In addition to the female mortality, male mortality was improved in large-scale mines, which may reflect the wealth effects of the large mining area. The marginal effects for sex ratio and average household size are also greater for the large-scale mines.
Table D.6: Additional Results: Heterogeneity with Respect to the Scale of Mines
Panel A: 1930 sample
Dependent Variable
(1) Population (2) Fertility (3) MR (Male) (4) MR (Female) (5) SR (6) Size
MineDeposit ($\leq$ median) 0.1790.179***0.179 ∗ ∗ ∗ 6.331-6.331*- 6.331 ∗ $-0.406$ $-0.466$ 0.0390.039***0.039 ∗ ∗ ∗ $0.008$
$(0.063)$ $(3.837)$ $(0.640)$ $(0.540)$ $(0.008)$ $(0.053)$
MineDeposit ($>$ median) 0.6580.658***0.658 ∗ ∗ ∗ 11.950-11.950***- 11.950 ∗ ∗ ∗ $-0.030$ $-0.524$ 0.0660.066***0.066 ∗ ∗ ∗ 0.221-0.221***- 0.221 ∗ ∗ ∗
$(0.080)$ $(3.322)$ $(0.483)$ $(0.413)$ $(0.012)$ $(0.044)$
FEs and Controls Yes Yes Yes Yes Yes Yes
Observations 1140 1140 1140 1140 1140 1140
Estimator OLS OLS OLS OLS OLS OLS
Panel B: 1935 sample
Dependent Variable
(1) Population (2) Fertility (3) MR (Male) (4) MR (Female) (5) SR (6) Size
MineDeposit ($\leq$ median) 0.2090.209***0.209 ∗ ∗ ∗ $2.179$ $-0.217$ $-0.338$ 0.0360.036***0.036 ∗ ∗ ∗ $-0.094$
$(0.072)$ $(3.357)$ $(0.480)$ $(0.599)$ $(0.009)$ $(0.064)$
MineDeposit ($>$ median) 0.6210.621***0.621 ∗ ∗ ∗ 8.586-8.586***- 8.586 ∗ ∗ ∗ 0.895-0.895**- 0.895 ∗ ∗ 0.915-0.915**- 0.915 ∗ ∗ 0.0650.065***0.065 ∗ ∗ ∗ $-0.058$
$(0.072)$ $(2.771)$ $(0.416)$ $(0.348)$ $(0.008)$ $(0.048)$
FEs and Controls Yes Yes Yes Yes Yes Yes
Observations 1364 1364 1364 1364 1364 1364
Estimator OLS OLS OLS OLS OLS OLS
***, **, and * represent statistical significance at the 1%, 5%, and 10% levels, respectively. Standard errors based on the heteroskedasticity-robust covariance matrix estimator are reported in parentheses.
Notes: MineDeposit ($\leq$ median) is an indicator variable for the municipalities within 5km from a mine, which had the number of miners less than or equal to the median. MineDeposit ($>$ median) is an indicator variable for the municipalities within 5km from a mine, which had the number of miners more than the median. Panels A and B show the results for the 1930 and 1935 samples, respectively. Column 1 shows the results for the log-transformed population. Column 2 shows the results for the marital fertility rate. Columns 3 and 4 show the results for the male and female mortality rates, respectively. Columns 5 and 6 show the results for the sex ratio and average household size, respectively.
References used in the Appendices
  • (1)
  • (2) Alberto Abadie. Semiparametric Difference-in-Differences estimators. The Review of Economic Studies, 72(1):1–19, January 2005.
  • (3)
  • (4) Takeshi Abe. Long-term recession in the interwar period and overcoming it. In Matao Miyamoto, editor, Japanese Economic History [in Japanese], pages 98–111. Hosōdaigaku kyōiku shinkōkai, Tokyo, 2008.
  • (5)
  • (6) Hunt Allcott and Daniel Keniston. Dutch disease or agglomeration? The local economic effects of natural resource booms in modern America. The Review of Economic Studies, 85(2):695–731, April 2018.
  • (7)
  • (8) Cecilia Nwadiuto Amadi, Zelinjo Nkeiruka Igweze, and Orish Ebere Orisakwe. Heavy metals in miscarriages and stillbirths in developing nations. Middle East Fertility Society Journal, 22(2):91–100, June 2017.
  • (9)
  • (10) Joshua D. Angrist, Guido W. Imbens, and Donald B. Rubin. Identification of Causal Effects using Instrumental Variables. Journal of the American Statistical Association, 91(434):444–455, June 1996.
  • (11)
  • (12) Fernando M Aragón and Juan Pablo Rud. Natural Resources and Local Communities: Evidence from a Peruvian Gold Mine. American Economic Journal: Economic Policy, 5(2):1–25, May 2013.
  • (13)
  • (14) Fernando M Aragón and Juan Pablo Rud. Polluting industries and agricultural productivity: Evidence from mining in Ghana. The Economic Journal, 126(597):1980–2011, September 2015.
  • (15)
  • (16) Manuel Arellano. Computing standard errors for robust within-groups estimators. Oxford Bulletin of Economics and Statistics, 49(4):431–434, November 1987.
  • (17)
  • (18) Anja Benshaul-Tolonen. Local Industrial Shocks and Infant Mortality. The Economic Journal, 129(620): 1561–1592, May 2019.
  • (19)
  • (20) Brian Beach and W Walker Hanlon. Coal Smoke and Mortality in an Early Industrial Economy. The Economic Journal, 128(615): 2652–2675, November 2017.
  • (21)
  • (22) Dan Black, Terra McKinnish, and Seth Sanders. The economic impact of the coal boom and bust. The Economic Journal, 115(503): 449–476, April 2005.
  • (23)
  • (24) Tuvia Blumenthal. Japanese economy in interwar period: International comparisons. In Takafusa Nakamura, editor, An Analysis of Japanese Economy in Interwar Period [in Japanese], pages 31–51. Tokyo, 1981.
  • (25)
  • (26) Francesco Caselli and Wilbur John Coleman II. The U.S. structural transformation and regional convergence: A reinterpretation. The Journal of Political Economy, 109(3): 584–616, June 2001.
  • (27)
  • (28) Clément de Chaisemartin and Xavier D’Haultfœuille. Two-way fixed effects and differences-in-differences with heterogeneous treatment effects: a survey. The Econometrics Journal, page utac017, June 2022.
  • (29)
  • (30) Fabian F Drixler. Hidden in Plain Sight: Stillbirths and Infanticides in Imperial Japan. The Journal of Economic History, 76(3):651–696, September 2016.
  • (31)
  • (32) Saburo Fumoto. Mitsubishi iizukaō shi (History of Mitsubishi Iizuka Coal Mine). [in Japanese]. Mitsubishi Mining Company, Tokyo, 1961.
  • (33)
  • (34) Christian B Hansen. Asymptotic properties of a robust variance matrix estimator for panel data when is large. Journal of Econometrics, 141(2): 597–620, December 2007.
  • (35)
  • (36) Guido W. Imbens, and Joshua D. Angrist. Identification and Estimation of Local Average Treatment Effects. Econometrica, 62(2):467–475, March 1994.
  • (37)
  • (38) Matao Miyamoto. Modern economic growth. In Matao Miyamoto, editor, Japanese Economic History [in Japanese], pages 53–111. Hosōdaigaku kyōiku shinkōkai, Tokyo, 2008.
  • (39)
  • (40) Mikio Sumiya. Nihon sekitan sangyō bunseki (Study of Japanese Coal Industry). [in Japanese]. Iwanami Shoten, Tokyo, 1968.
  • (41)
  • (42) Takumi Nagao. Kyūshū ko daisanki sō no sōjyo (Paleogene stratigraphy of Kyūshū). [in Japanese]. Nagasaki Earth Science Association, Nagasaki, 1975.
  • (43)
  • (44) Yutaka Nishinarita. Jyoshirōdō no shoruikei to sono henyō. In Masanori Nakamura, editor, Technological Innovation and Women’s Labor [in Japanese], pages 71–106. Tokyo Daigaku Shuppan Kai, Tokyo, 1985.
  • (45)
  • (46) Yoshihiro Ogino. Chikuhōtankō rōshi kankeishi (History of Labor-Management Relations in Chikuhō Coal Mine). [in Japanese]. Kyushi University Press, Fukuoka, 1993.
  • (47)
  • (48) Christian Siegel. Female relative wages, household specialization and fertility. Review of Economic Dynamics, 24: 152–174, 2017.
  • (49)
  • (50) Jean-Philippe C Stijns. Natural resource abundance and economic growth revisited. Resources Policy, 30(2): 107–130, June 2005.
  • (51)
  • (52) Naoki Tanaka and Yoshihiro Ogino. Miners protection issue and rationalization of coal extraction mechanism [in Japanese]. Enerugī-shi kenkyū, 8: 74–82, June 1977.
  • (53)
  • (54) Naoki Tanaka. Kindainihon tankōrōdōshi kenkyū (Research on the History of Coal Mining Labor in Modern Japan). [in Japanese]. Sōfūkan, Tokyo, 1984.
  • (55)
  • (56) James Tobin. Estimation of relationships for limited dependent variables. Econometrica, 26(1): 24–36, January 1958.
  • (57)
Documents used in the Appendices
  • (1)
  • (2) Aiikukai. Shussan, shusshō, shizan oyobi nyūjisibō tōkei (The statistics of birth, live births, fetal deaths, and infant deaths). [in Japanese]. Aiikukai, Tokyo, 1935.
  • (3)
  • (4) Association for Harmonious Cooperation. Zenkoku kōjyōkōzan meibo (The lists of factories and mines, 1932 edition). [in Japanese]. Association for Harmonious Cooperation, Tokyo, 1932.
  • (5)
  • (6) Association for Harmonious Cooperation. Zenkoku kōjyōkōzan meibo (The lists of factories and mines, 1937 edition). [in Japanese]. Association for Harmonious Cooperation, Tokyo, 1937.
  • (7)
  • (8) Compilation Committee of the Chronology of Chikuhō Coal Mining Industry. Chikuhōtankō kōgyōshi nenpyō (Chronology of Chikuhō Coal Mininig Industry). [in Japanese]. West Japan Cultural Association, Fukuoka, 1973.
  • (9)
  • (10) Fukuoka Prefecture. Fukuokaken tōkeisho, shōwa gonen daisanhen (Fukuoka Prefectural Statistics, volume 3 of 1930 edition). [in Japanese]. Fukuoka Prefecture, Fukuoka, 1932.
  • (11)
  • (12) Japan Mining Association. Nihon kōzan kyōkai siryō, dai isshū (Japan Mining Association Materials, volume 1). [in Japanese]. Japan Mining Association, Tokyo, 1928.
  • (13)
  • (14) Municipal Autonomy Research Association. Zenkoku shichōsonmei hensen sōran (General Survey of Changes in the Names of Municipalities). [in Japanese]. Nihon Kajo Publishing Co., Tokyo, 2006.
  • (15)
  • (16) Statistics Bureau of the Cabinet. Jinkōdōtaitōkei (The vital statistics of Japan, 1926–1940 editions). [in Japanese]. Statistics Bureau of the Cabinet, Tokyo, 1926–1942.
  • (17)
  • (18) Statistics Bureau of the Cabinet. Rodōtōkei jitchi chōsa hōkoku, Taishō jyūsan nen, Kōzan no bu (Labor statistics field survey report, 1924 edition, Mine section). [in Japanese]. Statistics Bureau of the Cabinet, Tokyo, 1926.
  • (19)
  • (20) Statistics Bureau of the Cabinet. Shichōsonbetsu jinkōdōtaitōkei (The vital statistics for municipalities, 1930 edition). [in Japanese]. Statistics Bureau of the Cabinet, Tokyo, 1932.
  • (21)
  • (22) Statistics Bureau of the Cabinet. Shōwagonen Kokuseichōsahōkoku, fukenhen (1930 Population Census of Japan, prefecture part). [in Japanese]. Statistics Bureau of the Cabinet, Tokyo, 1933
  • (23)
  • (24) Statistics Bureau of the Cabinet. Shōwagonen Kokuseichōsahōkoku, fukenhen (1930 Population Census of Japan, prefectural part). [in Japanese]. Statistics Bureau of the Cabinet, Tokyo, 1935a.
  • (25)
  • (26) Statistics Bureau of the Cabinet. Shōwagonen Kokuseichōsahōkoku, daiikkan (1930 Population Census of Japan, volume 1). [in Japanese]. Statistics Bureau of the Cabinet, Tokyo, 1935b.
  • (27)
  • (28) Statistics Bureau of the Cabinet. Shichōsonbetsu jinkōdōtaitōkei (The vital statistics for municipalities, 1935 edition). [in Japanese]. Statistics Bureau of the Cabinet, Tokyo, 1938.
  • (29)
  • (30) Statistics Bureau of the Cabinet. Shōwajyun̄en Kokuseichōsahōkoku, daiikkan (1935 Population Census of Japan, volume 1). [in Japanese]. Statistics Bureau of the Cabinet, Tokyo, 1939.
  • (31)
  • (32) Social Welfare Bureau. Shussan, shusshō, shizan oyobi nyūyōjisibō tōkei (The statistics of birth, live births, fetal deaths, and infant and child deaths). [in Japanese]. Social Welfare Bureau, Tokyo, 1941.
  • (33)